### v1

# A Classification Theorem for Steady Euler Flows

Abstract
Fix an analytic simply connected domain $\Omega\subset\mathbb{R}^{2}.$ A smooth function $\psi:\Omega\rightarrow\mathbb{R}$ that is constant on $\partial\Omega$ is said to define a steady Euler solution on $\Omega$ if the Poisson bracket between $\psi$ and its Laplacian vanishes. In other words, $\psi$ is a steady Euler solution if
$\{\psi,\Delta\psi\}=0.$ ($\star$)
[…]
1 Introduction
The study of steady Euler flows is old and vast, the first such solutions appearing in Euler’s original paper [19]. The interest in steady flows is natural, both from a mathematical and practical point of view. From the mathematical side, steady flows play a distinguished role in the long-time dynamics of all solutions as they do in any dynamical system. Indeed, the first step to understand the long-time dynamics of any system is to have a deep and complete understanding of the steady solutions. While it is well known that the Euler equation admits an infinite-dimensional family of steady states on any two-dimensional domain, little is known about the global properties of the set of solutions. The purpose of this work is to take a step in that direction.
1.1 The Euler equation
We begin by recalling the 2d Euler equation for incompressible flow on a simply connected domain $\Omega\subset\mathbb{R}^{2}$:
$\partial_{t}\omega+u\cdot\nabla\omega=0,$ (1.1)
$u=\nabla^{\perp}\Delta^{-1}\omega.$ (1.2)
Here, $\omega$ is the vorticity of the fluid and $u$ is its velocity field. The symbol $\nabla^{\perp}$ refers to the rotated gradient $\nabla^{\perp}=(-\partial_{2},\partial_{1})$, while $\Delta^{-1}$ is the inverse of the Dirichlet Laplacian on $\Omega.$ It is a well-known fact that sufficiently smooth solutions to (1.1)-(1.2) enjoy global existence and uniqueness, though $\omega$ may lose its smoothness in the long-time limit, $t\rightarrow\infty.$
One of the most captivating aspects of the 2d Euler equation is in the long-time behavior of solutions. An amazing fact, borne out of numerical and physical experiments, is that 2d Euler solutions starting from “any old data” tend to relax to much simpler states in the long-time limit. This is conjectured to occur, despite the conservative nature of the equation and the absence of a clear relaxation mechanism. Recent breakthroughs have been achieved in our understanding of this in certain perturbative regimes starting with the work [4] and then the later works [27, 28, 35]. See also [15] for a recent work leveraging the classical stability theory of steady states to make some long-time statements. Understanding the mechanism for this type of behavior outside of the perturbative regime appears well outside of the reach of any of the available techniques in the field. In order to be able to study such large data phenomena, perhaps in the long-time limit, it is necessary to build a flexible theoretical framework that will allow us to study questions of a more global nature. For us, this begins with having a robust understanding of Euler steady states.
1.2 Steady States
When studying steady solutions, it is helpful to introduce the stream function
$\psi=\Delta^{-1}\omega,$
in which case the steady Euler equation can be written as:
$\nabla^{\perp}\psi\cdot\nabla\Delta\psi:=\{\psi,\Delta\psi\}=0,$ (1.3)
where we note that $\psi$ is constant on $\partial\Omega.$ Classically, there are three classes of solutions that have been considered:
  • Solutions for which $\psi$ is invariant under translation in a fixed direction $e$, called shear flows.
  • Solutions for which $\psi$ is invariant under rotations about a fixed point $x_{0},$ called radial flows.
  • Solutions for which $\psi$ satisfies a semilinear elliptic equation:
    $\Delta\psi=F(\psi),$ (1.4)
    for some function $F:\mathbb{R}\rightarrow\mathbb{R}.$
The presence of the first two is clearly due to the symmetries of the Laplacian and the Poisson bracket. The third type is based on the simple observation that
$\{\psi,F(\psi)\}=F^{\prime}(\psi)\{\psi,\psi\}=0,$
for any $F\in C^{1}.$ Interestingly, it was only recently shown that there exist steady states strictly [^1] It is easy to find $C^{\infty}$ solutions consisting of two compactly supported radial bumps with respect to different centers that technically aren’t of any of these types, though they are locally radial. outside of these three types in the work [20] with some regularity and with any finite regularity in the work [18]. Both of these works rely on highly non-trivial bifurcation methods.
These three classes of steady solutions have a very nice structure. A quite robust linear stability theory has been developed for the first two, through the study of the Rayleigh equation [13, 31, 41]. The stability theory for the third type can be investigated using properties of the linearized operator associated to (1.4):
$\mathcal{L}=\Delta-F^{\prime}(\psi),$
see [1, 2, 3, 32, 33]. Given the tools the community has developed over the years to study these three types of steady states, it would be quite fortunate if we were able to show that these are the only steady states out there. This, and more, has been established locally in some non-degenerate settings [9, 10, 11, 17] (here, non-degenerate means that the Schrödinger operator $\mathcal{L}$ above is invertible).
In recent years, there has appeared a quite large literature on rigidity results for steady states. These results can be classified into two types: those making structural assumptions and those making smallness assumptions. In a series of papers, Hamel and Nadirashvili [23, 24, 25] established a number of rigidity results on symmetric domains assuming, for example, that the velocity field does not vanish on the domain. For example, they proved in [25] that sufficiently regular steady states on a regular annulus, say $B_{2}(0)\setminus B_{1}(0)$, with non-vanishing velocity must be radial. Of course, as is exemplified by the Dirichlet eigenfunctions, the non-vanishing of the velocity field is essential to such a result. See also the work of Li, Lv, Shahgholian, and Xie [30] on rigidity of Poiseuille flows in a strip and the work of Gui, Xie, and Xu[22] that studies the rigidity of steady Euler flows in terms of the flow angles. Another interesting recent result, among others, given by Gómez-Serrano, Yao, Park, and Shi [21] establishes the radial symmetry of solutions with compactly supported and non-negative vorticity. On the side of rigidity results under smallness conditions, there are now a few results showing that any solution sufficiently close to some special shear flow must be a shear flow [11, 34]. As an example, Lin and Zeng [34] proved that traveling-wave Euler solutions that are sufficiently close to the Couette flow on $\mathbb{T}\times[0,1]$ must be steady shear flows. Finally, we mention the work of Constantin, Drivas, and Ginsberg [10] that establishes results on both flexibility and rigidity in general domains, the main question being whether stable solutions inherit the symmetries of the domain. See also the works [7, 8, 16, 37] for further results on flexibility and rigidity in various contexts. We now state our main result.
1.3 Main Result
Theorem 1.1.
Let $\Omega$ be an analytic, simply connected, and bounded subset of $\mathbb{R}^{2}$. Assume that $\psi$ is analytic on $\bar{\Omega}$, is constant on $\partial\Omega$, and satisfies the steady Euler equation. Let $[a,b]$ be the range of $\psi$ on $\bar{\Omega}.$ Then, one of the following must hold.
  • $\Omega$ is a ball and $\psi$ is radial.
  • There exists $F\in C([a,b])$ that is analytic on the open interval $(a,b)$, for which
    $\Delta\psi=F(\psi).$
Remark 1.2.
We show moreover that $F$ is actually Hölder continuous globally, admitting a convergent Puiseux series at the endpoints. In fact, the precise form of the Puiseux expansion of $F$ at $a$ and $b$, see Lemma 4.5, allows us to define $\mathcal{L}=\Delta-F^{\prime}(\psi)$ as a self-adjoint operator on suitable (weighted) spaces. In particular, it can be decomposed into an invertible Schrödinger operator plus a compact perturbation and the Fredholm alternative can be applied to study its invertibility properties. This is key to understanding the local structure of the set of steady states near arbitrary ones. We summarise this as an interesting consequence of the proof in the following corollary.
Corollary 1.3.
Assume $\psi\in C^{\omega}(\bar{\Omega})$ for an analytic, simply connected, and bounded domain $\Omega.$ If
$\Delta\psi=F(\psi),$
and if $[a,b]$ is the range of $\psi$ on $\bar{\Omega},$ then $F$ is analytic on $(a,b)$ and admits a (locally) convergent Puiseux series at the endpoints.
The main novelty of Theorem 1.1 compared to other rigidity results on steady Euler flows mentioned above is that we make no structural or smallness assumptions on $\Omega$ or on $\psi$. The results are also, in some sense, optimal since [18] shows that we cannot relax the analyticity condition to finite smoothness and Theorem 4.12 shows the result cannot hold in general on multiply connected domains. The analyticity assumption is one drawback of Theorem 1.1; it would be nice if it could be replaced by a less stringent structural assumption.
1.4 Main ideas
We now move to discuss some of the main ideas that go into the proof of Theorem 1.1. As we have mentioned above, the proof of Theorem 1.1 is based on a number of observations about the Poisson bracket, the zero sets of analytic functions, and about over-determined elliptic problems. To clarify the main points of the argument, we will briefly sketch some of the main ideas here.
1.4.1 Some observations about $\{f,g\}=0$
It is well known, using a straightening lemma, that if $f,g\in C^{1}$ and $\{f,g\}=0$ in a neighborhood of some $x_{*}\in\mathbb{R}^{2}$, while $\nabla f(x_{*})\not=0,$ then we can write
$g=F(f),$
for some $F\in C^{1}$ in a sufficiently small neighborhood of $x_{*}.$ This is a simple application of the inverse function theorem. It is also clear that this need not be the case globally. There are two important examples one should keep in mind in this regard, both of which we give on $\mathbb{T}^{2}$ (though the examples can be easily modified to work on the unit disk, for example).
Example 1: Non-Analytic $g$. Take $f(x,y)=\sin(x)\sin(y).$ In this case, each regular level set of $f$ has two connected components consisting of simple closed curves. We may take $g$ to simply be a bump function that is constant along the level sets of $f$ restricted to the first quadrant and also compactly supported inside the first quadrant. In this case, $g$ cannot be expressed as a function of $f$ globally on $\mathbb{T}^{2}$ (since $f$ takes the same values in the first and third quadrants, but $g$ takes different values there). We give two illustrative figures as follows.
Figure 1: 2D Contour plot of $g$
Figure 2: 2D Contour plot of $f$
Example 2: A wall of critical points. Take $f=\sin^{2}(x)\sin^{2}(y)$ and $g=\sin(x)\sin(y).$ Then, $g$ cannot be written as $F(f).$ In this case, $f$ and $g$ are both analytic; however, the set of critical points of $f$ partitions the domain into multiple cells surrounded by walls of critical points. We give two illustrative figures below. The red lines in the figure below are the set of critical points of $f$.
Figure 3: 2D contour plot of $g$
Figure 4: 2D contour plot of $f$
It turns out that these examples encapsulate the only two obstructions to a global functional relation between $f$ and $g.$ We now state a simple consequence of our key proposition in this regard.
Lemma 1.4.
Assume that $f,g$ are analytic functions on some open, connected, domain $\Omega.$ Assume that $\{f,g\}=0$ and that the critical points of $f$ of even order are isolated. Then, there exists a continuous function $F:\mathbb{R}\rightarrow\mathbb{R}$ so that
$g=F(f).$
In particular, in order that there does not exist a global $F$, $f$ must have “walls” of critical points of even order (See Definition 2.1 for the precise meaning of critical point of even order).
1.4.2 Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0.$
Given the preceding discussion, the obstruction to having a global relation between $\Delta\psi$ and $\psi$ in a 2d Euler steady state is the existence of a closed curve of critical points of $\psi$ of even order. Since $\psi$ is analytic, we can consider the inner-most such curve, $\Gamma$. On the region enclosed by $\Gamma$, called $\Omega_{\Gamma}$, $\psi$ will satisfy a semi-linear elliptic equation with non-linearity $F$ and the full gradient of $\psi$ vanishes on $\Gamma$. Note that the curve $\Gamma$ may have singularities. Now there appear three cases:
  1. (i)
    $\psi$ vanishes higher than quadratically on $\Gamma$,
  2. (ii)
    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ and the associated non-linearity $F$ is not analytic at $c=\psi(\Gamma)$,
  3. (iii)
    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ with $F$ analytic at $c=\psi(\Gamma).$
In case (iii), it turns out that $F$ may be extended to a larger region containing $\Gamma$ and eventually we return to cases (i) and (ii). In both cases (i) and (ii), because $F$ is necessarily non-analytic at $c=\psi(\Gamma)$, it turns out that $\psi-c$ is necessarily single signed in $\Omega_{\Gamma}$–this is a key aspect of the argument. In case (i), then we can apply a known result of Brock [5] to deduce that $\psi$ must be radial; a nice feature of Brock’s result [5] is that it does not require the boundary of the domain to be smooth. The analysis in case (ii) is more difficult and will be discussed in more detail below. We give an illustration depicting the potential complexity of the level-set in the figure below [^2] The solid curves on the figure above are the set of critical points of $\psi$ in an elliptical domain. The boundary of region A demonstrates situation (i), where all the points are critical points of degree higher than 2. The boundary of region B demonstrates the second case (ii), in the case where all of the points are quadratic. Regions C and D could represent the third case, where $F$ could be extended from C to D, while D falls into the case (i) or (ii)..
Figure 5: Illustrations for a potential critical set of $\psi$
One snag in case (ii) is that we have to show that this curve $\Gamma$ is actually smooth and that $\psi$ is single signed in its interior. For general analytic functions, it is well known that the zero set may be quite complicated and may contain multiple singular points of different types [29, 6], even more so in high dimension. Note, for example, that the set of critical points in Example 2 above has a singularity at the origin. In the figure above, we have to show that case (ii) above actually corresponds to a region like B in the figure (i.e. a region with a smooth boundary). A key observation here is that, when we have that $\{\psi,\Delta\psi\}=0,$ the degree of vanishing of $\nabla\psi$ on each connected component of $\Gamma$ will necessarily be constant. This is established by using the constancy of $\Delta\psi$ on the level sets of $\psi$. Once the constancy of the degree is established, we use an elementary fact: singularities in the zero set of an analytic function occur only at points where the degree of vanishing of the function changes. The constancy of the degree of vanishing thus implies that $\Gamma$ is smooth.
1.4.3 Moving Plane Arguments
From the preceding two steps, we see that if a steady state $\psi$ does not satisfy the global relation $\Delta\psi=F(\psi)$, up to subtracting a constant, we end up with a simply connected domain $\Omega_{\Gamma}$ with analytic boundary where
$\Delta\psi=F(\psi)\quad\text{in}\,\,\Omega_{\Gamma},$
$\psi=\nabla\psi=0\quad\text{on}\,\,\partial\Omega_{\Gamma}.$
We may also assume that $F(s)\approx\sqrt{s}$ as $s\rightarrow 0.$ As mentioned above, it is not difficult to show that $\psi$ cannot change sign in $\Omega_{\Gamma}$. By scaling, we can assume that $0\leq\psi\leq 1$ Now, the overdetermined problem above seems similar to those considered by Serrin in his landmark paper [39]. Unfortunately, a key difference here is that we do not have the Lipschitz smoothness of the non-linearity $F,$ which is very important for the moving plane method used in [39]. Indeed, as we have already said, near $s=0,$ $F(s)\approx\sqrt{s}$; moreover, $F$ need not be smooth as $s\rightarrow 1$ (the maximum value of $\psi$) either. To deal with the singularity as $s\rightarrow 0,$ we use the fact that we have a lower bound on the growth of $\psi$ in the direction normal to the boundary, since the vorticity is non-zero (and constant) on the boundary. Then we invoke a refinement of the Hopf lemma for elliptic problems on small domains under some growth conditions on the coefficients. This effectively deals with the singularity as $s\rightarrow 0$. Remarkably, it turns out that the potential singularity as $s\rightarrow 1$ always has the right sign for the maximum principle (see the proof of Lemma 4.11).
1.5 Organization of the Paper
In Section 2, we establish some general results about analytic functions satisfying $\{f,g\}=0$ and recall a few results about the zero sets of analytic functions. The main consequence of this is that, if an analytic Euler solution $\psi$ does not satisfy an equation of the type (1.4), we are ensured the existence of a level set of $\psi$ on which the gradient vanishes. In Section 3, we prove a few results about these sets, using more fundamentally that $\psi$ is an Euler solution. In particular, we show the analyticity of the non-isolated portions of the critical points with quadratic vanishing. This sets us up with an overdetermined elliptic problem in a subdomain of the original domain. At the end of that section, we then establish the rigidity of solutions with curves of critical points with higher than quadratic vanishing. In Section 4, we establish the rigidity of solutions with a suitably chosen curve of critical points with quadratic vanishing first by deducing certain regularity properties of the function $F$ (defined in a suitable subdomain) and then by applying the moving plane method. In Section 4.3 we give a counterexample to Theorem 1.1 in an analytic non-simply connected bounded domain, hence showing the optimality of the statement.
2 Facts about $\{f,g\}=0$
The purpose of this section is to establish Proposition 2.2 as well as to recall some basic facts about the zero sets of analytic functions.
2.1 Key Facts
We start by giving a notation for the degree of flatness of an analytic function at a point.
Definition 2.1.
For any non-constant analytic $f:\Omega\rightarrow\mathbb{R},$ and $x_{*}\in\Omega,$ we define $d_{f}(x_{*})$ to be the first $k\geq 1$ for which $D^{k}(f)(x_{*})\not=0.$
We then introduce notation for the set of critical points of $f$ and those with even vanishing:
$\mathcal{C}(f)=\{x\in\Omega:d_{f}(x)>1\},\qquad\mathcal{C}_{e}(f)=\{x\in\Omega:d_{f}(x)\in 2\mathbb{N}\},$
and state the key proposition.
Proposition 2.2.
Assume that $\{f,g\}=0$ in some open connected set $\Omega\subset\mathbb{R}^{2}.$ Suppose that $\Gamma$ is an analytic curve connecting $x,y\in\Omega$ and that $\Gamma\cap\mathcal{C}_{e}(f)=\emptyset.$ Then, if $f(x)=f(y),$ it must be that $g(x)=g(y).$
To prove this proposition, let us first establish a generalization of the well-known straightening lemma.
Lemma 2.3.
Assume that $f,g$ are analytic and $\{f,g\}=0$ in a neighborhood of a point $x_{*}\in\Omega.$ Assume that $d:=d_{f}(x_{*})\not\in 2\mathbb{N}.$ Then, there exists a continuous function $F$ so that $g=F(f)$ in a neighborhood of $x_{*}.$ In fact, $F(s)$ is analytic in the variable $(s-f(x_{*}))^{1/d}$.
Remark 2.4.
The case $d=1$, which is that $\nabla f(x_{*})\not=0,$ is the standard straightening lemma which is valid only under the assumptions that $f$ and $g$ are $C^{1}$ in a neighborhood of $x_{*}.$
Remark 2.5.
Strictly speaking, we do not need the generalization to $d>1$ for the results proven later; however, it is conceptually important in order to understand the difference between “extendible” and “unextendible” domains later. The result may also be of independent interest. It should be noted that, in this generalization, analyticity plays a key role with the statement not holding true in the smooth class.
Proof.
Without loss of generality, by applying suitable translations and a rotation, set $x_{*}=0$, $f(x_{*})=0,$ and $\partial_{1}^{d}f(0)>0.$ Assume that $d$ is odd. The point is that we have that $\partial_{1}f(t,0)>0$ for $t\in[-\epsilon,0)\cup(0,\epsilon]$ for some $\epsilon>0$. In particular, the function $f(t,0)$ is real-analytic and invertible in $(-\epsilon,\epsilon)$. It follows that we can express its inverse using a Puiseux series for $t$ sufficiently small. In particular, there exists (a unique) $F$ that is analytic in the variable $s^{1/d}$ for which
$g(t,0)=F(f(t,0)),$
for every $t\in(-\frac{\epsilon}{2},\frac{\epsilon}{2}).$ Let $\kappa$ be the radius of convergence of the Puiseux series of $F$ at $0$ We now take $\delta<\frac{\epsilon}{2}$ be so small that $f(B_{\delta}(0))\subset(-\kappa,\kappa).$ We will show that
$g(x)=F(f(x)),$
for every $x\in B_{\delta}(0).$ Fix any $x\in B_{\delta}(0).$ First, take $x_{1}\in(-\delta,\delta)\setminus\{0\}$. By assumption, $(x_{1},0)$ is not a critical point of $f$. It follows that there is a real analytic $G$ so that $g(x)=G(f(x)),$ for every $x$ in a neighborhood of $(x_{1},0).$ In particular, there is an $\epsilon(x_{1})>0$ such that for all $t\in(x_{1}-\epsilon(x_{1}),x_{1}+\epsilon(x_{1})),$ we have that
$G(f(t,0))=F(f(t,0)).$
However, $f$ is monotone in that region. Therefore, $G=F$ in a neighborhood of $f(x_{1},0).$ By analyticity, this means that $F\equiv G$ on their domains.
At this point, we have shown that there is a ball around each point of the segment $\{(t,0):t\in(-\epsilon,\epsilon)\setminus\{0\}\}$ where $g=F(f).$ Now we will extend this to the whole ball $B_{\delta}(0).$ Fix any $x\in B_{\delta}(0)$ with $x_{1}\neq 0$. Without loss of generality, assume $x_{2}>0.$ Now consider the function $g(x_{1},t),$ where $t\in[0,x_{2}].$ This is an analytic function of $t$, since $g$ is an analytic. Now, $F(f(x_{1},t))$ is analytic and equal to $g$ in a neighborhood of $t=0$. Moreover, $F(f(x_{1},t))$ is also analytic except at the possibly discrete set of points where $f(x_{1},t)=0$, where $F(f(x_{1},t))$ has a convergent Puiseux series (note that $f(x_{1},0)\not=0$). It follows that $F(f(x_{1},t))$ is actually analytic on $[0,x_{2}]$ and that $g(x_{1},t)=F(f(x_{1},t))$ for every $t\in[0,x_{2}].$ In particular, $g(x_{1},x_{2})=F(f(x_{1},x_{2})).$
We are now in a position to prove Proposition 2.2.
Proof of Proposition 2.2.
For any point $x\in\Gamma$, by applying Lemma 2.3 to $f$ and $g$, we can find a neighborhood $U_{x}$ and a Puiseux function $F_{x}$ such that $g=F(f)$ in $U_{x}$. As the collection of the open set $U_{x}$ forms an open cover for the compact set $\Gamma$, we can choose $U_{x^{1}},\cdots,U_{x^{m}}$, such that $U_{x^{i}}\cap U_{x^{i+1}}\neq\emptyset$ and $\Gamma\subset\cup_{i=1}^{m}U_{x^{i}}$. In the case where $f$ is constant in $\Gamma$, as the pair $(g,f)$ satisfies a function relationship in each $U_{x^{i}}$, we have $g$ is constant in $U_{x^{i}}\cap\Gamma$. Notice that $\{U_{x^{i}}\}$ is a finite open cover for $\Gamma$, we have $g$ is constant in $\Gamma$.
In the case where $f$ is not constant in $\Gamma$, we claim $F_{x^{1}}$ can be extended to be a Puiseux function $F$ in the range of $f([\Gamma])$ with $g=F(f)$ in $\Gamma$. We first prove that $F_{x^{1}}$ can be extended to a Puiseux function $F$ in the range $f(U_{x^{1}}\cap\Gamma)\cup f(U_{x^{2}}\cap\Gamma)$, with $g=F(f)$ in $\Gamma\cap(U_{x^{1}}\cup U_{x^{2}})$. Here $F_{x^{1}}$ and $F_{x^{2}}$ are two Puiseux functions respectively in the range of $f(U_{x^{1}}\cap\Gamma)$ and $f(U_{x^{2}}\cap\Gamma)$, and the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$. As $f$ is not constant in $\Gamma$, the analytic assumption of $f$ and $\Gamma$ implies that $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$ contains an open interval in $\mathbb{R}$. Now due to the fact that $f(\Gamma\cap U_{x^{1}})$ and $f(\Gamma\cap U_{x^{2}})$ are connected, the condition that $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(U_{x^{1}}\cap U_{x^{2}})$ implies that the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}})\cap f(\Gamma\cap U_{x^{2}})$. Hence, we can extend $F_{x^{1}}$ with the aid of $F_{x^{2}}$ to a Puiseux function $F$ in $f(U_{x^{1}})\cup f(U_{x^{1}})$ and $g=F(f)$ in $\Gamma\cap(U_{x}^{1}\cup U_{x}^{2})$. We can iterate the above procedure to extend $F_{x^{1}}$ to a Puiseux function $F$ in the range of $f([\Gamma])$ such that $g=F(f)$ in $\Gamma$ and we now finish the proof. ∎
2.2 Zero Sets of Analytic Functions
In this section, we will briefly discuss the structure of zero set of a real analytic function in two variables. We firstly recall a classical result concerning the structure of zero set of real analytic function.
Proposition 2.6 (See Corollary 2 in [40]).
Let $f$ be an analytic function near $x^{0}$ and $f(x^{0})=0$, then one of the following three things must happen.
  1. 1.
    $x^{0}$ is an isolated point in the zero set for $f$.
  2. 2.
    $f\equiv 0$.
  3. 3.
    The nodal set $\{f=0\}$ near $x^{0}$ is a collection of an even number of analytic curves branching at $x^{0}$.
It is an immediate corollary of Proposition 2.6 that for any given point $x$, which is not isolated in the nodal set for a non-constant analytic function $f$, there are at least two analytic curves $\Gamma_{1}$, $\Gamma_{2}$ branching at $x$ and consisting of nodal points of $f$. In most cases, we cannot glue $\Gamma_{1}$ and $\Gamma_{2}$ together to form a single analytic curve. For example, in the case $f(x,y)=x^{2}-y^{3}$, the nodal set has a cusp at the origin. However, in the case where $\Gamma_{1}$ consists of critical points of the same degree (in particular, if the degree doesn’t jump up at the common point), $\Gamma_{1}\cup\Gamma_{2}$ is an analytic curve.
Lemma 2.7.
Let $x^{0}$ be a fixed point and we assume there is an analytic curve $\Gamma_{1}:[0,1]\rightarrow\mathbb{R}^{2}$ with $\Gamma_{1}(0)=x^{0}$, and $d_{f}(x)=m$ for all $x\in\Gamma_{1}([0,1])$. Then the level set $\{x|f(x)=f(x^{0})\}$ near $x^{0}$ is an analytic curve $\Gamma$. Moreover, for any point $x\in\Gamma$, $d_{f}(x)=m$.
Proof.
Without loss of generality, by applying suitable rotation, we assume $x^{0}=0$, $f(0)=0$ and $\partial_{1}^{m}f(0)>0$. By applying the analytic inverse function theorem to $\partial_{1}^{m-1}f$ at 00, we have the set $\{\partial_{1}^{m-1}f=0\}$ near 00 is an analytic curve $\Gamma$ passing through 00. The existence of the analytic curve $\Gamma_{1}$ consisting of critical points of degree $m$ that starts at 00 ensures that for infinitely many points $x^{i}$ accumulate at 00 in $\Gamma$, we have
$D^{k}f(x^{i})=0,\text{for $1\leq k\leq m-1$ and $i$. }$ (2.1)
As $\Gamma$ and $f$ are analytic, we have
$D^{k}f(x)=0,\text{for $x\in\Gamma$, and $1\leq k\leq m-1$.}$ (2.2)
Then due to the fundamental theorem of calculus, we have $f=0$ in $\Gamma$.
Now it suffices to find an open set $\mathcal{O}$ near the origin such that $\Gamma\cap\mathcal{O}$ consists of critical point of degree $m$ and $\{f=0\}\cap\mathcal{O}=\Gamma\cap\mathcal{O}$. We notice the condition $\partial_{1}^{m}f(0)>0$ ensures that $(1,0)$ is not a tangent vector of $\Gamma$ at 00. Now as $f$ and $\Gamma$ are analytic, we can choose $\mathcal{O}:=(-\epsilon,\epsilon)^{2}$ to be a small cube near 00 with the requirement that all the points $x$ in $\mathcal{O}$ satisfy $\partial_{1}^{m}f>0$ and in addition, for any point $x\in\mathcal{O}\setminus\Gamma$, there is a horizontal line $l_{x}$ connecting $x$ and a point $x_{\Gamma}$ in $\Gamma$. Now as $\partial_{1}f(x_{\Gamma})=\cdots=\partial_{1}^{m-1}f(x_{\Gamma})=0$ and $\partial_{1}^{m}f>0$ in $l_{x}$, by applying mean value theorem to $f$ in $l_{x}$, we have $f(x)\neq 0$. ∎
3 Structure of the set of critical points
From here and on, we fix a single analytic steady state $\psi,$ satisfying (1.3) on $\bar{\Omega}.$ Towards proving Theorem 1.1, we will also assume that there is no $F$ for which $\psi$ solves (1.4) on all of $\Omega.$ By Proposition 2.2, as well as Proposition 2.6, it must thus be that $\psi$ has curves of critical points. The purpose of this section is to study these in more detail. To start, the equation
$\{\psi,\Delta\psi\}=0$
implies that $\Delta\psi$ is constant on any connected level set of $\psi.$ We thus have the following crucial Lemma.
Lemma 3.1.
Let $\Gamma$ be a curve of points in $\mathcal{C}.$ Suppose that for some $x_{*}\in\Gamma$, $d_{\psi}(x_{*})=2.$ Then, $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $\Gamma.$
Proof.
By Proposition 2.6, for a non-isolated critical point $x$, we have that $D^{2}\psi(x)=0$ is equivalent to $\Delta\psi=0$; moreover, $\Delta\psi$ is constant along any curve. This is because, if we take any analytic curve emanating from $x,$ and denote by $\tau$ and $n$ the tangent and normal vectors to the curve at $x,$ we have that $\partial_{\tau n}\psi(x)=\partial_{\tau\tau}\psi(x)=0.$
Using Lemma 2.7, we get the following corollary.
Corollary 3.2.
The points for which $d_{\psi}(x)=2$ consist of a collection of isolated points or isolated analytic closed loops in $\Omega$.
The following result, which will be used later, is of a similar flavor.
Lemma 3.3.
Let $\Gamma$ be an analytic curve of points of $\psi$ of degree $d\geq 3$. Then, for all $x\in\Gamma,$
$d_{\Delta\psi}(x)=d-2.$
Proof.
Fix $x\in\Gamma$. By the rotational invariance of $\Delta,$ we may assume that the tangent vector to $\Gamma$ at $x$ is $(1,0).$ Now, for each $y\in\Gamma$ we have $D^{d-1}\psi(y)=0$, from which we conclude that $D^{d-1}\partial_{1}\psi(x)=0$. It thus follows that $D^{d}\psi(x)\not=0$ is equivalent to $\partial_{2}^{d}\psi(x)\not=0.$ This, in turn, is equivalent to $\partial_{2}^{d-2}\Delta\psi(x)\not=0.$
Remark 3.4.
We note that it is actually possible to show that the degree of vanishing of an analytic Euler steady state has to be constant on any connected component of the set of critical points. This is shown in the thesis of the second named author [26].
3.1 The Inner-Most Loop
Since $\mathcal{C}(\psi)$ contains curves, $\Omega$ must be partitioned by the level set $\{\nabla\psi=0\}$ (this follows from the above discussion and Proposition 2.6). In particular, we can find at least one curve of critical points that is either a closed loop in $\Omega$ or starts and ends at $\partial\Omega.$ Let $\Gamma$ be one such curve with the property that the area it encloses contains no curves of critical points. Let us call this enclosed domain $\Omega_{\Gamma},$ which is necessarily simply connected. All of the analysis below will be restricted to $\Omega_{\Gamma}.$ Since $\Omega_{\Gamma}$ only contains isolated critical points of $\psi,$ we have that
$\Delta\psi=F(\psi)\text{ in $\overline{\Omega_{\Gamma}},$ }$ (3.1)
for some continuous function $F.$ By subtracting a constant, we may also assume that
$\psi|_{\partial\Omega_{\Gamma}}=0.$
We will now show that we have, in fact, that $\nabla\psi|_{\partial\Omega_{\Gamma}}=0,$ which is automatic when $\Gamma$ is a loop.
Lemma 3.5.
If $\Gamma$ touches $\partial\Omega,$ then the portion of $\partial\Omega$ between the two endpoints of $\Gamma$ lies in $\mathcal{C}(\psi).$
Proof.
By Corollary 3.2, if $\Gamma$ touches $\partial\Omega$, then $d_{\psi}(x)\geq 3$ for all $x\in\Gamma.$ Moreover, $\psi$ must then vanish on $\partial\Omega.$ Now, take some $x\in\partial\Omega.$ If $\nabla\psi(x)\not=0,$ then $F$ from (3.1) must be analytic in a neighborhood of $0$ In addition, the condition that $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$ shows $F(0)=0$. This, however, contradicts Lemma 3.3, since $F$ being analytic with $F(0)=0$ would imply that $\Delta\psi=F(\psi)$ vanishes to the same order (or higher) than $\psi$ on $\Gamma.$
Corollary 3.6.
One of the following two can hold
  • Case 1: $\partial\Omega_{\Gamma}$ is analytic and contains only critical points of degree 2.
  • Case 2: $\partial\Omega_{\Gamma}$ is a closed loop containing critical points of degree larger than 2.
Our goal will now be to show that, in both cases, $\psi$ is necessarily a radial function. We will first deal with Case 2, which will be solved essentially by applying existing results from the literature. Case 1 will be dealt with in the following section using the moving plane method.
3.2 Rigidity in Case 2
In the previous subsection, we reduced the problem into studying an semi-linear elliptic equation of Serrin type [39]. More precisely, we found a simply connected domain $\Omega_{\Gamma}$ for which $\nabla\psi$ vanishes on $\partial\Omega_{\Gamma}$ and for which $\Omega_{\Gamma}$ contains only isolated critical points of $\psi$. As a consequence of Lemma 1.4, we deduced that $\psi$ solves (3.1) in $\Omega_{\Gamma}$. In this section, we consider Case 2 of Corollary 3.6, where all the points of $\partial\Omega_{\Gamma}$ are critical points with degree greater than or equal to $3$. Our goal will be to establish the following proposition.
Proposition 3.7.
Let $\Omega_{\Gamma}$ be a simply connected domain and $\psi$ be an analytic function in $\overline{\Omega_{\Gamma}}$ such that there is no non-isolated critical points of $\psi$ in $\Omega_{\Gamma}\cap\{\psi=0\}$. Now we further assume there is a continuous function $F$, such that $\psi$ satisfies the following over-determined elliptic system.
$\Delta\psi(x)=F(\psi)(x),\text{ for $x\in\Omega_{\Gamma}$,}$ (3.2)
$\psi(x)=\nabla\psi(x)=\nabla^{2}\psi(x)=0,\text{ for all $x\in\partial\Omega_{\Gamma}$},$
then $\psi$ is a radial function.
We first recall the following result based on Steiner symmetrization method.
Proposition 3.8 (From Theorem 5.2 and Theorem 6.1 in [5] or Theorem 3.2 in [38]).
Let $\Omega$ be an open bounded set and $\psi$ a non-negative $C^{2}$ solution to
[…]
Then $F$ being continuous implies that $\psi$ is locally radial that is:
$\mathbb{D}:=\{0<\psi<sup_{x\in\Omega}\psi(x)\}=A\cup\mathcal{C},$
where $A=\cup_{i}A_{i}$, $\mathcal{C}=\{\nabla\psi=0\}$ and $A_{k}$ are disjoint open annuli or balls and $\psi$ is radially symmetric in $A_{k}$.
Proof of Proposition 3.7.
Given Proposition 3.8, all we need to do is show that $\psi$ is single signed on $\Omega_{\Gamma}.$ Towards a contradiction, assume that $\psi$ takes both positive and negative values in $\Omega_{\Gamma}.$ Simply by continuity, it follows that the set of zeros of $\psi$ in $\Omega_{\Gamma}$ is infinite. It follows, by our choice of $\Gamma,$ that at least one of those zeros is not a critical point of $\psi$ (since the critical set in $\Omega_{\Gamma}\cap\{\psi=0\}$ is discrete). It follows that $F$ is analytic in a neighborhood of zero. This contradicts Lemma 3.3 (just like in the proof of Lemma 3.5). ∎
4 Moving Plane Argument
We now move on to prove the rigidity in Case 1 of Corollary 3.6. In particular, we may assume that there exists a simply connected domain $\Omega_{\Gamma}$ with $\partial\Omega_{\Gamma}$ analytic with $\psi$ satisfying
$\Delta\psi(x)=F(\psi)(x),\text{for $x\in\Omega_{\Gamma}$,}$ (4.1)
$\psi(x)=\nabla\psi(x)=0,\text{for all $x\in\partial\Omega_{\Gamma}$}.$
Moreover, by scaling, we may assume that $F(0)=1.$ In particular,
$\partial_{nn}\psi|_{\partial\Omega_{\Gamma}}=1.$
An important consequence is that $\psi$ is increasing in the direction normal to the boundary, in a tubular neighborhood of the boundary. In the case where $F$ is $C^{1}$ and $\psi$ is non-negative, we have the following rigidity result due to Serrin [39] (see also the nice review [36]).
Theorem 4.1 (Serrin, 1971).
Assume that $\psi$ satisfies (4.1) and that $\psi>0$ in $\Omega_{\Gamma},$ while $F$ is Lipschitz continuous. Then, $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
In our case, we cannot assume that $F$ is $C^{1}$ in general. In fact, there are two cases.
Lemma 4.2.
It is either that
  • $F(s)$ is not analytic at $s=0,$ in which case $\psi$ is non-negative on $\Omega_{\Gamma}.$ This is called the unextendable case.
  • $F(s)$ is analytic at $s=0.$ This is called the extendable case.
4.1 The Unextendable Case
This is the case where $F(s)$ is not analytic in $s$ at 00. As mentioned in Lemma 4.2, $\psi$ must be non-negative in $\Omega_{\Gamma}$, since otherwise there would be infinitely many points where $\psi=0$ in $\Omega_{\Gamma}$. Without loss of generality, by scaling and multiplication by a constant factor, we assume that $\max_{\Omega_{\Gamma}}\psi=1$. Analogously as in the proof of Proposition 3.7, the condition that $\psi$ does not have a non-isolated critical point in $\Omega_{\Gamma}$ implies that $F$ is analytic in $(0,1)$. Our goal in this section is to prove the following proposition:
Proposition 4.3.
Let $\Omega_{\Gamma}$ be a simply connected analytic domain, $\psi$ be a non-negative analytic function in $\overline{\Omega_{\Gamma}}$. Assume $F\in C([0,1])$ is analytic on $(0,1).$ If $F$ is not analytic at 00 and $\psi$ satisfies the over-determined elliptic equation (4.1), then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
Remark 4.4.
Without a loss of generality, we may assume $F(0)=1,$ since the case $F(0)=0$ is already covered by Proposition 3.7.
For the proof of Proposition 4.3, it suffices to prove that $\psi$ is even symmetric to all directions $e\in\mathbb{S}^{1}$. Without loss of generality, by rotation, $e=(1,0)$. We will use the moving plane method to show that $\psi$ is even symmetric to $(1,0)$. We start with a series of lemmas concerning the regularity of $F$ and the critical set of $\psi$. First, either $F$ is analytic at $1$ or $F$ is not smooth and admits a Puiseux expression near $1$.
Lemma 4.5.
When the maxima of $\psi$ are critical points of degree $2k_{0},$ for some $k_{0}\geq 1,$ $F(s)$ admits the following convergent expansion.
$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{ for $0\leq 1-s\ll 1$,with $a_{k}$ real and $a_{k_{0}-1}<0$. }$ (4.2)
Remark 4.6.
The sign condition $a_{k_{0}-1}<0$ will play an important role in the moving plane argument.
Proof.
Fix a point $x^{0}\in\Omega_{\Gamma}$ such that $\psi(x^{0})=1$, as $\psi$ attains the global maximum at $x^{0}$, we have $d_{\psi}(x^{0})$ is an even number. In the case $d_{\psi}(x^{0})=2$, without loss of generality, by a suitable rotation and translation, we assume $x^{0}=0$ and $\partial_{11}\psi(0)<0$. As $\partial_{1}\psi(0)=0$ and $\partial_{11}\psi(0)<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$ and $f_{-}(s)$, with $f_{+}(s)\geq 0$ and $f_{-}(s)\leq 0$. We show that $f_{+}$ and $f_{-}$ admit a Puisuex expression near $1$. In particular, we will show that there is an analytic function $G$, such that $f_{+}(s)=G(\sqrt{1-s})$ and $f_{-}(s)=G(-\sqrt{1-s})$. Indeed given the Taylor expression of $\psi(t,0)=1+a_{2}t^{2}+\sum_{n\geq 3}a_{n}t^{n}$ with $a_{2}<0$, we have
$\psi(f_{\pm},0)=s\Leftrightarrow f_{\pm}^{2}\left(1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}\right)=\frac{s-1}{a_{2}}\Leftrightarrow f_{\pm}\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}}=\pm\sqrt{\frac{s-1}{a_{2}}}.$ (4.3)
Now the function $s\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}s^{k-2}}$ is an analytic function with a non-vanishing derivative at 00, hence it admits an analytic inverse $H$ and we can correspondingly choose $G(s)=H(\frac{s}{\sqrt{-a_{2}}})$. With the aid of $f_{+}$, $f_{-}$ and $G$ in the claim, for $0\leq 1-s\ll 1$, we have
$F(s)=\Delta\psi(G(\sqrt{1-s}),0)=\Delta\psi(G(-\sqrt{1-s}),0).$
The fact that $\Delta\psi(G(\sqrt{1-s}),0)$ agrees with $\Delta\psi(G(-\sqrt{1-s}),0)$ shows that $F$ is analytic near $1$.
Now in the case where $d_{\psi}(x^{0})=2k_{0}\geq 4$, again assume that $x^{0}=0$ and $\partial_{1}^{2k_{0}}\psi<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$, $f_{-}(s)$. Similarly as in the case where $d_{\psi}(x^{0})=2$, we can show that $F$ near $1$ has the Puiseux expression.
$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{with $a_{k}$ real, $a_{k_{0}-1}<0$, for $0\leq 1-s\ll 1$. }$ (4.4)
Analogously, by taking a normal line segment through a boundary point $x\in\partial\Omega_{\Gamma}$, we have $F(s)$ is analytic in the variable $\sqrt{s}$ near 00. Together with the condition that $F$ is not analytic near 00, we have the following Puiseux expression of $F$ near 00.
Lemma 4.7.
$F$ near 00 admits a convergent Puiseux series expression:
$F(s)=\sum_{k=0}a_{k}s^{\frac{k}{2}},\text{for $0\leq s\ll 1$ , with $a_{k}$ real, $a_{0}=1$ and $a_{k_{0}}\neq 0$ for an odd number $k_{0}$.}$ (4.5)
Moreover, by arguing similarly as in the case $d_{\psi}(x^{0})=2$ in the proof of Lemma 4.5, the condition $F$ is not analytic at 00 implies that $\psi$ is positive in $\Omega_{\Gamma}$. Together with fact
$\partial_{nn}\psi|_{\partial\Omega_{\Gamma}}=1,$
we can show $\psi$ has a uniform lower bound in $\Omega_{\Gamma}$.
Lemma 4.8.
There exists $C>0$ such that for all $x\in\Omega_{\Gamma}$,
$\frac{dist(x,\partial\Omega_{\Gamma})^{2}}{C}\leq\psi(x)\leq Cdist(x,\partial\Omega_{\Gamma})^{2}.$ (4.6)
Now we are ready to apply the moving plane method to show that $\psi$ is even symmetric in the direction $e=(1,0)$. Without loss of generality, by translation, we fix $\inf_{x\in\Omega_{\Gamma}}x_{1}=0$. For $\lambda>0$, let $H_{\lambda}:=\{x_{1}=\lambda\}$. We now reflect $\psi$ and $\Omega_{\Gamma}$ through the hyperplane $H_{\lambda}$:
$\pi_{\lambda}(x_{1},x_{2}):=(2\lambda-x_{1},x_{2}),\ \psi_{\lambda}(x):=\psi(\pi_{\lambda}(x)),\ \Omega_{\lambda}:=\{x_{1}<\lambda\}\cap\Omega_{\Gamma},\text{ and }\breve{\Omega}_{\lambda}:=\pi_{\lambda}(\Omega_{\lambda}).$
We are going to show $\psi_{\lambda}$ is always less than or equal to $\psi$ in $\breve{\Omega}_{\lambda}$ before $\partial\breve{\Omega}_{\lambda}$ touches the physical boundary $\partial\Omega_{\Gamma}$ tangentially. We first define the reflection monotonicity property for $\psi$ at level $\lambda$.
Definition 4.9.
We say $\psi$ has the reflection monotonicity property at the level $\lambda$, if $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$ and $\psi\geq\psi_{\lambda}$ in $\breve{\Omega}_{\lambda}$.
Since $\partial\Omega_{\Gamma}$ is analytic, for $\lambda\ll 1$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. Moreover, as $\nabla\psi=0$ and $\partial_{nn}\psi=1$ in $\Omega_{\Gamma}$, $\psi$ satisfies the reflection monotonicity property at the level $0<\lambda\ll 1$ We will now prove that the reflection monotonicity property persists unless $\partial\breve{\Omega}_{\lambda}$ hits $\partial\Omega_{\Gamma}$ tangentially.
Lemma 4.10.
Define
$\lambda_{0}=\sup\{\lambda|\text{$\psi$ satisfies the reflection monotonicity property at level $\lambda$}\},$
we have $h_{\lambda_{0}}=\psi-\psi_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$. Moreover, at least one of the following occurs when $\lambda=\lambda_{0}$.
  • Internal tangency. $\partial\breve{\Omega}_{\lambda_{0}}$ is tangent to $\partial\Omega_{\Gamma}$ at some point $x^{0}\notin H_{\lambda_{0}}$.
  • Boundary tangency. $\partial\breve{\Omega}_{\lambda_{0}}$ is tangent to $\partial\Omega_{\Gamma}$ at a point $x^{0}\in H_{\lambda_{0}}$. In particular, the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is $(0,1)$.
Figure 6: Internal tangency
Figure 7: Boundary tangency
The proof of Lemma 4.10 is based on the observation that as long as $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$, the difference $\psi-\psi_{\lambda}$ satisfies a singular elliptic equation where the singular coefficent has a uniform upper bound.
Lemma 4.11.
For all $\lambda>0$, such that $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma},$ $h_{\lambda}:=\psi-\psi_{\lambda}$ satisfies a singular elliptic equation
$\Delta h_{\lambda}=c_{\lambda}(x)h_{\lambda},$ (4.7)
where the singular coefficient $c(x)$ has a uniform upper bound
$c_{\lambda}^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}.$ (4.8)
Proof.
Due to the reflection symmetry of the semilinear elliptic equation, $\Delta\psi_{\lambda}=F(\psi_{\lambda})$, and thus
$\Delta h_{\lambda}=\frac{F(\psi)-F(\psi_{\lambda})}{\psi-\psi_{\lambda}}h_{\lambda}:=c_{\lambda}(x)h_{\lambda}.$ (4.9)
In the case where $F$ is analytic at $1$, as $F$ is already analytic in $(0,1)$, from Lemma 4.7, there exists a function $G\in C^{1}([0,1])$ and a constant $d$ such that $F(s)=d\sqrt{s}+G(s)$ ($d$ might be zero, which corresponds to the case where the Puiseux expression of $F$ at 00 has no $\sqrt{s}$ term). Now we have
$c_{\lambda}=\frac{d}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}$ (4.10)
Due to the lower bound of $\psi$ in the equation (4.6), we have
$\psi_{\lambda}>\frac{dist(x,\partial\breve{\Omega}_{\lambda})^{2}}{C},$ (4.11)
together with the fact that $\psi$ and $\psi_{\lambda}$ enjoy a uniform $C^{1}$ upper bound, we have $|c_{\lambda}|\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, where $M>0$ is independent of $\lambda$.
In the case where $F$ is not analytic near $1$, since $F$ is analytic in $(0,1)$, Lemma 4.5 and Lemma 4.7 show that there exists an integer $k_{0}\geq 2$, a negative constant $d_{1}<0$, a constant $d_{2}$ and a function $G\in C^{1}([0,1])$, such that $F(s)=d_{1}(1-s)^{\frac{k_{0}-1}{k_{0}}}+d_{2}\sqrt{s}+G(s)$. We now have the corresponding expression for $c_{\lambda}$
$c_{\lambda}=d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}+\frac{d_{2}}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}.$ (4.12)
As $d_{1}<0$, the term $d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}$ is non-negative. Hence, similiarly as in the case where $F$ is analytic at $1$, we have $c_{\lambda}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, for a positive $M$ independent of $\lambda$. ∎
We now prove Lemma 4.10, arguing by contradiction.
Proof of Lemma 4.10.
Due to the continuity of $\psi$, $h_{\lambda_{0}}:=\psi-\psi_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$. Now we assume that at the level of $\lambda_{0}$, none of the internal tangency or boundary tangency occurs, in particular, there is a $\delta^{*}>0$, such that for $\lambda_{0}\leq\lambda<\lambda_{0}+\delta^{*}$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. We will show that there is a $\delta\in(0,\delta^{*})$ such that $\psi$ satisfies the reflection monotonicity at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$. We start by proving that $h_{\lambda_{0}}$ is positive in $\breve{\Omega}_{\lambda_{0}}$. By Lemma 4.11, $h_{\lambda_{0}}$ is a nonnegative solution to $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ with $c^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$. Now suppose that there is a point $x^{0}\in\breve{\Omega}$ such that $h_{\lambda}(x^{0})=0$, we automatically have $\nabla h_{\lambda}(x^{0})=0$. Then the Proposition A.1 in the appendix shows in a small ball whose boundary contains $x^{0}$, $h_{\lambda_{0}}$ is identically zero. Due to the analytic regularity of $\psi$, it further implies that $\psi_{\lambda_{0}}=\psi$. Hence, $\Omega_{\Gamma}$ is even symmetric to $H_{\lambda_{0}}$ and we get a contradiction. Hence, we now have
$h_{\lambda_{0}}(x)>0,\text{ for all $x\in\breve{\Omega}_{\lambda_{0}}$.}$ (4.13)
Anaglously, for any point $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$, we may apply Proposition A.1 to a small ball that hits $x$ boundary-tangentially at $x$, and conclude that
$\partial_{1}h_{\lambda_{0}}(x)>0,\text{ for all $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$},$ (4.14)
unless $h_{\lambda_{0}}\equiv 0.$
In the rest of the proof, we will show that there is a $\delta>0$ such that $\psi$ satisfies the reflection monotonicity property at level $\lambda\in[\lambda_{0},\lambda+\delta].$ First, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit $\partial\Omega_{\Gamma}$ boundary-tangentially, for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta^{*}]$, $H_{\lambda}\cap\partial\Omega_{\Gamma}$ consists of two points $x_{\lambda}^{1}$, $x_{\lambda}^{2}$($x_{\lambda}^{1}$ is above $x_{\lambda}^{2}$). Moreover, there is a $\epsilon_{0}>0$ such that $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}}^{1})<-\epsilon_{0}$ and $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}})(x_{\lambda_{0}}^{2})>\epsilon_{0}$. As $\partial_{12}h_{\lambda}(x_{\lambda}^{1})$ and $\partial_{12}h_{\lambda}(x_{\lambda}^{2})$ are uniformly continuous in $\lambda$ , we choose $\delta_{1}>0$ such that
$\partial_{12}h_{\lambda}(x_{\lambda}^{1})<-\frac{1}{2}\epsilon_{0},\partial_{12}h_{\lambda}(x_{\lambda}^{2})>\frac{1}{2}\epsilon_{0},\text{ for$\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta_{1}$}.$ (4.15)
As $x_{\lambda}^{1}$ is a critical point of $\psi$ and $\psi_{\lambda}$, we have
$\nabla h_{\lambda}(x_{\lambda}^{1})=0,$ (4.16)
moreover by symmetry, we have
$\partial_{11}h_{\lambda}(x_{\lambda}^{1})=\partial_{22}h_{\lambda}(x_{\lambda}^{1})=0.$ (4.17)
Due to the smoothness of $\partial\Omega_{\Gamma}$ and that the boundary tangency doesn’t occur when $\lambda=\lambda_{0}$, there exists a $\epsilon>0$, such that the direction $(1,-\epsilon)$ is always pointing to the exterior of $\breve{\Omega}_{\lambda}$ at $x_{\lambda}^{1}$ for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$(we may decrease the value of $\delta_{1}$ when it is necessary). Now since $D^{3}h_{\lambda}$ is uniformly bounded, from the equations (4.15), (4.16) and (4.17), there exists $r_{1}>0$, such that $h_{\lambda}>0$ in $B_{r_{1}}(x_{\lambda}^{1})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$. Analogously, in $B_{r_{1}}(x_{\lambda}^{2})$ (we decrease the value of $\delta_{1}$ when necessary), $h_{\lambda}\geq 0$ in $B_{r_{1}}(x_{\lambda}^{2})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$. Now due to the equation (4.14), there is a $\epsilon_{1}>0$ such that we have $\partial_{1}h_{\lambda_{0}}(x)>\epsilon_{1}$ for all $x\in H_{\lambda_{0}}\setminus(B_{r_{1}}(x_{\lambda_{0}}^{1})\cup B_{r_{1}}(x_{\lambda_{0}}^{2}))$. Similar to the previous case, we can find a $r_{2}>0$ and a $\delta_{2}\in(0,\delta_{1})$, such that $h_{\lambda}\geq 0$ in the set $T_{\lambda}:\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r_{2}\}\setminus(B_{r}(x_{\lambda}^{1})\cup B_{r}(x_{\lambda}^{2}))\}$ for all $\lambda_{0}\leq\lambda\leq\lambda+\delta_{2}$.
In summary, we show there is a uniform $r>0$ and $\delta>0$, such that if $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$, we have $h_{\lambda}\geq 0$ in $N_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r\}$, which is a tubular neighborhood of $H_{\lambda}$. Note by (4.13) and Lemma 4.8, there is a $\epsilon_{2}>0$, such that $h_{\lambda_{0}}>\epsilon_{2}$ on the set $\breve{\Omega}_{\lambda_{0}}\setminus N_{\lambda_{0}}$. Now by the uniform continuity of $h_{\lambda}$, we may decrease the value of $\delta$ (still make it positive), such that $h_{\lambda}>\frac{\epsilon_{2}}{2}$ on the set $\breve{\Omega}_{\lambda}\setminus N_{\lambda}$ for all $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$. ∎
Now we apply Hopf’s lemma to prove that $\psi$ and $\Omega_{\Gamma}$ are even symmetric to the $(1,0)$ direction to finish the proof of Proposition 4.3.
Proof.
Lemma 4.10 shows there exists $\lambda_{0}>0$ with $h_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$ and at least one of the internal boundary tangency and boundary tangency happens for $\breve{\Omega}_{\lambda_{0}}$. In the case where internal tangency occurs at $x^{0}\in\partial\Omega_{\Gamma}$, by Lemma 4.11, $h_{\lambda_{0}}$ is a non-negative solution of the equation $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ in $\breve{\Omega}_{\lambda_{0}}$, with $c^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$. Moreover, we have $\nabla h_{\lambda_{0}}(x^{0})=\nabla\psi(x^{0})-\nabla\psi_{\lambda_{0}}(x^{0})=0$. Now we apply Proposition A.1 for a sufficiently small ball $B\subset\breve{\Omega}_{\lambda_{0}}$ that hits $\breve{\Omega}_{\lambda_{0}}$ internal-tangentially at $x^{0}$, and conclude $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$.
Now we consider the case where the boundary of tangency happens at $x^{0}\in H_{\lambda_{0}}\cap\partial\Omega_{\Gamma}$. First, similar to above, we have $\nabla h_{\lambda_{0}}(x^{0})=0$. In addition, $\partial\breve{\Omega}_{\lambda_{0}}$ forms a right angle at $x^{0}$, and the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is parallel to direction $(0,1)$ In particular, we have $\nabla^{2}h_{\lambda_{0}}(x^{0})=0$. Now analogously as in the case of the internal tangency, by applying Proposition A.2 in the appendix to a small half ball contained in $\breve{\Omega}_{\lambda_{0}}$ that hits $\Omega_{\Gamma}$ boundary-tangentially, we have $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$. ∎
4.2 The Extendable Case
Now we assume that $F$ is analytic at $s=0.$ Using the argument in the proof of Proposition 2.2, we see that we can extend the definition of $F$ so that $\psi$ satisfies
$\Delta\psi=F(\psi)$
on an extended domain $\tilde{\Omega}_{\Gamma}$, whose boundary is again either an analytic curve of points of degree 2, or a curve of points of degree 3 or higher. By the Jordan curve theorem, $\Gamma$ is either in the interior of the $\partial\Omega_{\Gamma}$ or in the exterior. In either case, since we are assuming that $F$ cannot be extended to all of $\Omega$, we will hit a curve of critical points along which $F$ is not analytic. In that case, we take the smallest curve inscribed in it with that property and we apply either Proposition 3.7 or 4.3 depending on the degree of vanishing of $\nabla\psi$ on the boundary. The proof of Theorem 1.1 is now complete.
4.3 Counterexample
It is important to note that the rigidity in Theorem 1.1 is optimal in terms of regularity in bounded domains. First, the results of [18] show that Theorem 1.1 does not hold for flows that have finite smoothness even for the weaker assumption of non-locally-radial flows. Secondly, Theorem 1.1 does not hold in the case where the compact domain $\Omega$ is not simply connected, as the following result shows.
Theorem 4.12.
Fix an analytic Jordan curve $\gamma\subset\mathbb{R}^{2}$. Then there exists a tubular neighborhood $\Omega$ of $\gamma$ and a stream function $\psi\in C^{\omega}(\bar{\Omega})$ constant on $\partial\Omega$ solving
$\begin{cases}\Delta\psi=2+\psi^{\frac{5}{2}}\text{ in }\Omega^{interior}\\ \Delta\psi=2-\psi^{\frac{5}{2}}\text{ in }\Omega^{exterior}\end{cases},\text{ and }\psi=\frac{\partial\psi}{\partial n}=0\text{ on }\gamma.$
Thus, any non-circular $\gamma$ gives rise to a steady Euler flow $\psi$ that is analytic in a neighborhood of $\gamma$ that is neither radial nor a solution of a semilinear elliptic equation of the form (1.4). The proof of Theorem 4.12 is inspired by the classical proof the Cauchy-Kovalevskaya Theorem and a careful application of the so-called majorization method to a well chosen majorization system. A detailed proof is given in the thesis of the second named author [26].
5 Conclusion and Future Directions
We gather here a few interesting directions and problems that can be considered, in view of Theorem 1.1 and its proof. We have already mentioned that replacing the analyticity assumption with some other less stringent assumption would be desirable. We mention some other interesting problems below.
5.1 Local and Global Questions
Given Theorem 1.1 and the extremely large literature on semilinear elliptic equations, an important problem is to study the local and global properties of the set of Euler steady states. As mentioned in the introduction, the invertibility properties of the linear Schrödinger operator
$\mathcal{L}=\Delta-F_{*}^{\prime}(\psi_{*})$
play an important role in the local structure of steady states (see [9, 10, 11, 12, 34]). One consequence of Theorem 1.1 and the proof is that we can always define $\mathcal{L}$ as a (unbounded) self-adjoint operator on $L^{2}$. This gives hope to have a general local understanding of the manifold of steady states, and then possibly a global one.
Most of the previous results on the local structure of steady states have assumed the invertibility of $\mathcal{L}$. It is a very interesting problem to study the general case where $\mathcal{L}$ has a kernel. From the classical Fredholm theory, we know the kernel is finite dimensional. When the kernel of $\mathcal{L}$ is one-dimensional, it is still possible to give a quite clear picture of the local structure of steady states (though this has not yet been spelled out in the literature). Once the kernel of $\mathcal{L}$ is two-dimensional or higher, the local structure of steady states may be quite complicated [11]. It is a very interesting problem to study this in general and this is related to the (im)possibility of the existence of isolated states [14]. If we had a good understanding of the local structure of steady states in the general case, we could hope to patch together this information to get more global information about the structure of the set of analytic Euler steady states. It is possible that some of these issues could be simpler to study on generic domains [42]. For example, on generic domains it may even be possible to establish some kind of global correspondence between $\psi$ and $F$ in the analytic class [9].
In general, it would be very interesting to have results related to the connectedness, path-connectedness, and smoothness of the manifold of steady states [12].
5.2 Rigidity questions on other domains
A first obvious question is whether similar results can be established in different domains. We have already shown that it does not hold on some annular domains, though it is not clear whether a modification holds on $\mathbb{T}\times[0,1],$ for example. A very interesting problem is to determine whether a version of Theorem 1.1 can hold on $\mathbb{T}^{2}.$ In that case, the question would be whether all analytic steady states are necessarily either a shear flow or a solution to a semilinear elliptic equation.
Acknowledgements
T.M.E. acknowledges partial funding from the NSF DMS-2043024, the Alfred P. Sloan foundation, and the Simons Foundation. He also acknowledges helpful conversations with T. Drivas and A. Kiselev. Y.H. acknowledges partial funding from NSF DMS-2043024. A. R. S. is supported by the ERC/UKRI advanced grant SWAT. The research of C. X. is partially supported by NSFC grant 12250710674, and Program of Shanghai Academic Research Leader 22XD1421400.
Appendix A Hopf Lemmas for elliptic equations with singular coefficents
In this section, we list some known facts about the maximum principle for the solutions to the elliptic equation $-\Delta h+c(x)h=0$ in a Lipschitz domain $\Omega$. Here $c(x)$ could be unbounded, but it satisfies the following uniform upperbound $c^{-}(x)\leq\frac{M}{dist(x,\partial\Omega)}$. The results we present here are crucial in our proof of Proposition 4.3.
Proposition A.1 (The Hopf Lemma in small balls, see Proposition 4.3 in [38]).
We consider a non-negative smooth solution $h$ to the following elliptic equation in the ball $B_{r}(0,r):$
$-\Delta h+c(x)h=0.$ (A.1)
Now, if we further assume $h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial h}{\partial x_{2}}(0)>0$.
Proposition A.2 (The Hopf Lemma in small half balls, see Proposition 4.4 in [38]).
Let $B_{r}^{+}:=B_{r}(0,-r)\cap\{x_{1}\geq 0\}$, we consider a non-negative smooth solution $h$ to the following elliptic equation in $B_{r}^{+}:$
$-\Delta h+c(x)h=0\text{ with $c^{-}(x)\leq\frac{M}{dist(x,0)}$.}$ (A.2)
Now if we further assume $\nabla h(0)=h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial^{2}h}{\partial\eta^{2}}(0)>0$, for any direction $\eta=(\eta_{1},\eta_{2})$ satisfying $\eta_{1}>0$, $\eta_{2}<0$ in $B_{r}^{+}$.

### v2

@@
 A Classification Theorem for Steady Euler Flows
-Abstract
-Fix an analytic simply connected domain $\Omega\subset\mathbb{R}^{2}.$ A smooth function $\psi:\Omega\rightarrow\mathbb{R}$ that is constant on $\partial\Omega$ is said to define a steady Euler solution on $\Omega$ if the Poisson bracket between $\psi$ and its Laplacian vanishes.
-In other words, $\psi$ is a steady Euler solution if
-$\{\psi,\Delta\psi\}=0.$ ($\star$)
-[…]
-[…]
-[…]
-1 Introduction
+Abstract.
+Fix an analytic simply connected domain $\Omega\subset\mathbb{R}^{2}.$ We show that all analytic steady states of the Euler equations with stream function $\psi$ are either radial or solve a semi-linear elliptic equation of the form $\Delta\psi=F(\psi)$ with Dirichlet boundary conditions.
+In particular, if $\Omega$ is not a ball, then there exists a one to one correspondence between analytic steady states of the Euler equations and analytic solutions of equations of the form $\Delta\psi=F(\psi)$ with Dirichlet boundary conditions.
+Key words and phrases:
+Steady Euler equations, stream function, analytic, critical point, moving plane method
+1991 Mathematics Subject Classification:
+35Q31, 35N25, 35J15, 35B53
+1. Introduction and Main Results
 The study of steady Euler flows is old and vast, the first such solutions appearing in Euler’s original paper [24].
-The interest in steady flows is natural, both from a mathematical and practical point of view.
-From the mathematical side, steady flows play a distinguished role in the long-time dynamics of all solutions as they do in any dynamical system.
-Indeed, the first step to understand the long-time dynamics of any system is to have a deep and complete understanding of the steady solutions.
-While it is well known that the Euler equation admits an infinite-dimensional family of steady states on any two-dimensional domain, little is known about the global properties of the set of solutions.
+The interest in steady flows is natural, both from a mathematical and a practical point of view.
+From the mathematical side, steady flows play a distinguished role in the long-time dynamics of all solutions, as they do in any dynamical system.
+Indeed, the first step to understanding the long-time dynamics of any system is to have a deep and complete understanding of the steady solutions.
+While it is well known that the Euler equations admit an infinite-dimensional family of steady states on any two-dimensional domain, little is known about the global properties of the set of solutions.
 The purpose of this work is to take a step in that direction.
-1.1 The Euler equation
-We begin by recalling the 2d Euler equation for incompressible flow on a simply connected domain $\Omega\subset\mathbb{R}^{2}$:
-$\partial_{t}\omega+u\cdot\nabla\omega=0,$ (1.1)
-$u=\nabla^{\perp}\Delta^{-1}\omega.$ (1.2)
+1.1. The Euler Equations
+We begin by recalling the 2D Euler equations for incompressible flow on a simply connected bounded domain $\Omega\subset\mathbb{R}^{2}$:
+(1.1) $\partial_{t}\omega+u\cdot\nabla\omega=0,$
+(1.2) $u=\nabla^{\perp}\Delta^{-1}\omega.$
 Here, $\omega$ is the vorticity of the fluid and $u$ is its velocity field.
@@
 This is conjectured to occur, despite the conservative nature of the equation and the absence of a clear relaxation mechanism.
-Recent breakthroughs have been achieved in our understanding of this in certain perturbative regimes starting with the work [4] and then the later works [27, 28, 35].
+Recent breakthroughs have been achieved in our understanding of this in certain perturbative regimes, starting with the work [4] and then the later works [32, 33, 40].
 See also [19] for a recent work leveraging the classical stability theory of steady states to make some long-time statements.
-Understanding the mechanism for this type of behavior outside of the perturbative regime appears well outside of the reach of any of the available techniques in the field.
+Understanding the mechanism for this type of behavior outside the perturbative regime appears to be well outside the reach of any of the available techniques in the field.
 In order to be able to study such large data phenomena, perhaps in the long-time limit, it is necessary to build a flexible theoretical framework that will allow us to study questions of a more global nature.
 For us, this begins with having a robust understanding of Euler steady states.
-1.2 Steady States
+1.2. Steady States
 When studying steady solutions, it is helpful to introduce the stream function
 $\psi=\Delta^{-1}\omega,$
-in which case the steady Euler equation can be written as:
-$\nabla^{\perp}\psi\cdot\nabla\Delta\psi:=\{\psi,\Delta\psi\}=0,$ (1.3)
-where we note that $\psi$ is constant on $\partial\Omega.$ Classically, there are three classes of solutions that have been considered:
+in which case the steady Euler equations can be written as:
+(1.3) $\nabla^{\perp}\psi\cdot\nabla\Delta\psi:=\{\psi,\Delta\psi\}=0,$
+where we note that $\psi$ is chosen to be zero on $\partial\Omega.$ Classically, there are three classes of solutions that have been considered:
   • Solutions for which $\psi$ is invariant under translation in a fixed direction $e$, called shear flows.
@@
   • Solutions for which $\psi$ satisfies a semilinear elliptic equation:
-    $\Delta\psi=F(\psi),$ (1.4)
+    (1.4) $\Delta\psi=F(\psi),$
     for some function $F:\mathbb{R}\rightarrow\mathbb{R}.$
-The presence of the first two is clearly due to the symmetries of the Laplacian and the Poisson bracket. The third type is based on the simple observation that
+The presence of the first two is due to the symmetries of the Laplacian and the Poisson bracket. The third type is based on the simple observation that
 $\{\psi,F(\psi)\}=F^{\prime}(\psi)\{\psi,\psi\}=0,$
@@
 A quite robust linear stability theory has been developed for the first two, through the study of the Rayleigh equation [18, 37, 48].
-The stability theory for the third type can be investigated using properties of the linearized operator associated to (1.4):
+The stability theory for the third type can be investigated using properties of the linearized operator associated with (1.4):
 $\mathcal{L}=\Delta-F^{\prime}(\psi),$
 see [3, 2, 1, 38, 39].
-Given the tools the community has developed over the years to study these three types of steady states, it would be quite fortunate if we were able to show that these are the only steady states out there.
+Given the tools the community has developed over the years to study these three types of steady states, it would be great if we were able to show that these are the only steady states out there.
 This, and more, has been established locally in some non-degenerate settings [12, 13, 15, 22] (here, non-degenerate means that the Schrödinger operator $\mathcal{L}$ above is invertible).
@@
 Of course, as is exemplified by the Dirichlet eigenfunctions, the non-vanishing of the velocity field is essential to such a result.
-See also the work of Li, Lv, Shahgholian, and Xie [30] on rigidity of Poiseuille flows in a strip and the work of Gui, Xie, and Xu[22] that studies the rigidity of steady Euler flows in terms of the flow angles.
-Another interesting recent result, among others, given by Gómez-Serrano, Yao, Park, and Shi [21] establishes the radial symmetry of solutions with compactly supported and non-negative vorticity.
-On the side of rigidity results under smallness conditions, there are now a few results showing that any solution sufficiently close to some special shear flow must be a shear flow [11, 34].
+See also the work of Li, Lv, Shahgholian, and Xie [35] on the rigidity of Poiseuille flows in a strip and the work of Gui, Xie, and Xu[27] that studies the rigidity of steady Euler flows in terms of flow angles.
+Another interesting recent result, among others, given by Gómez-Serrano, Park, Shi, and Yao [25] establishes the radial symmetry of solutions with compactly supported and non-negative vorticity.
+On the side of the results of rigidity under smallness conditions, there are now some results showing that any solution sufficiently close to some special shear flow must be a shear flow [15, 36].
 As an example, Lin and Zeng [36] proved that traveling-wave Euler solutions that are sufficiently close to the Couette flow on $\mathbb{T}\times[0,1]$ must be steady shear flows.
@@
 See also the works [8, 9, 21, 42] for further results on flexibility and rigidity in various contexts.
-We now state our main result.
-1.3 Main Result
+We now state our main results.
+1.3. Main Results
 Theorem 1.1.
-Let $\Omega$ be an analytic, simply connected, and bounded subset of $\mathbb{R}^{2}$.
-Assume that $\psi$ is analytic on $\bar{\Omega}$, is constant on $\partial\Omega$, and satisfies the steady Euler equation.
+Let $\Omega$ be a simply connected and bounded domain of $\mathbb{R}^{2}$.
+Assume that $\psi$ is analytic on $\bar{\Omega}$, is constant on $\partial\Omega$, and satisfies the steady Euler equation (1.3).
 Let $[a,b]$ be the range of $\psi$ on $\bar{\Omega}.$ Then, one of the following must hold.
@@
 Remark 1.1.
+As $\psi$ is analytic in $\overline{\Omega}$ and $\partial\Omega$ is a level set of $\psi$, due to the structure theorem on the zero sets of real analytic functions (See Proposition 2.3), the boundary $\partial\Omega$ has to be a piecewise analytic curve.
+Remark 1.2.
 We show moreover that $F$ is actually Hölder continuous globally, admitting a convergent Puiseux series at the endpoints.
 In fact, the precise form of the Puiseux expansion of $F$ at $a$ and $b$, see Lemma 4.4, allows us to define $\mathcal{L}=\Delta-F^{\prime}(\psi)$ as a self-adjoint operator on suitable (weighted) spaces.
-In particular, it can be decomposed into an invertible Schrödinger operator plus a compact perturbation and the Fredholm alternative can be applied to study its invertibility properties.
+In particular, it can be decomposed into an invertible Schrödinger operator plus a relatively compact perturbation.
+Hence the Fredholm theory can be applied to study its invertibility properties.
 This is key to understanding the local structure of the set of steady states near arbitrary ones.
-We summarise this as an interesting consequence of the proof in the following corollary.
+We summarize this as an interesting consequence of the proof in the following corollary.
 Corollary 1.2.
-Assume $\psi\in C^{\omega}(\bar{\Omega})$ for an analytic, simply connected, and bounded domain $\Omega.$ If
+For a given analytic, simply connected and bounded domain $\Omega$, Suppose that $\psi\in C^{\omega}(\bar{\Omega})$ (the set of real-analytic functions in $\bar{\Omega}$) takes a constant on $\partial\Omega$.
+If
 $\Delta\psi=F(\psi),$
-and if $[a,b]$ is the range of $\psi$ on $\bar{\Omega},$ then $F$ is analytic on $(a,b)$ and admits a (locally) convergent Puiseux series at the endpoints.
-The main novelty of Theorem 1.1 compared to other rigidity results on steady Euler flows mentioned above is that we make no structural or smallness assumptions on $\Omega$ or on $\psi$.
-The results are also, in some sense, optimal since [18] shows that we cannot relax the analyticity condition to finite smoothness and Theorem 4.12 shows the result cannot hold in general on multiply connected domains.
-The analyticity assumption is one drawback of Theorem 1.1; it would be nice if it could be replaced by a less stringent structural assumption.
-1.4 Main ideas
-We now move to discuss some of the main ideas that go into the proof of Theorem 1.1.
-As we have mentioned above, the proof of Theorem 1.1 is based on a number of observations about the Poisson bracket, the zero sets of analytic functions, and about over-determined elliptic problems.
+and if $[a,b]$ is the range of $\psi$ in $\bar{\Omega},$ then $F$ is analytic on $(a,b)$ and admits a (locally) convergent Puiseux series at the endpoints.
+The main novelty of Theorem 1.1, compared with other results of rigidity in steady Euler flows mentioned above, is that we make no structural or smallness assumptions on $\Omega$ or on $\psi$.
+The results are also, in some sense, optimal, since [23] shows that we cannot relax the analyticity condition to finite smoothness, and Proposition 4.11 shows that the result cannot hold in general on multiply connected domains.
+Remark 1.3.
+The analyticity of the velocity field is the major assumption of Theorem 1.1.
+It would be nice if one could give a necessary and sufficient condition for the regularity of the velocity field so that Theorem 1.1 holds.
+On the other hand, for any $s>1$, there exists a single-variable compactly supported function in the Gevrey class $G^{s}$ (see [34]).
+Hence, we can use two compactly supported disjoint radial bumps to construct a steady flow whose stream function belongs to the Gevrey class $G^{s}$ with $s>1$, but is neither radial nor a solution to a global semilinear elliptic equation.
+Therefore, in terms of regularity, analyticity might be the optimal assumption for velocity fields.
+Remark 1.4.
+The simple connectedness of the domain is necessary.
+In Section 4.3, a counterexample is provided to illustrate that in a multiply connected domain, there exists a flow that can be neither described by a solution of global semilinear equation for the stream function nor a radial flow.
+Remark 1.5.
+As an application of Theorem 1.1, we can show that in a simply connected convex bounded domain, any analytic steady state whose vorticity is single-signed should inherit the mirror symmetry of the domain.
+In particular, any analytic steady flow with distinguished-signed vorticity in the disk is a radial flow.
+For more details on this property, one may refer to the Appendix B.
+Remark 1.6.
+To better understand the long-time behavior of unsteady flow, it would be desirable to establish a similar result in a broader class of steady states.
+An important fact that could help us to get closer is that every steady state with $\psi\in C^{2,\alpha}$ has analytic Lagrangian trajectories [10, 45, 14].
+In fact, this can be extended to all steady states with merely bounded vorticity.
+The main idea is that steady Euler solutions with bounded vorticity actually have some stratified regularity away from the critical points, as in the case of regular vortex patches [44, 11, 5].
+We will present this result in a forthcoming paper.
+1.4. Main Ideas
+We now move on to discuss some of the main ideas that go into the proof of Theorem 1.1.
+As we have mentioned above, the proof of Theorem 1.1 is based on a number of observations about the Poisson bracket, the zero sets of analytic functions, and over-determined elliptic problems.
 To clarify the main points of the argument, we will briefly sketch some of the main ideas here.
-1.4.1 Some observations about $\{f,g\}=0$
-It is well known, using a straightening lemma, that if $f,g\in C^{1}$ and $\{f,g\}=0$ in a neighborhood of some $x_{*}\in\mathbb{R}^{2}$, while $\nabla f(x_{*})\not=0,$ then we can write
-$g=F(f),$
-for some $F\in C^{1}$ in a sufficiently small neighborhood of $x_{*}.$ This is a simple application of the inverse function theorem.
+1.4.1. Some observations about $\{f,g\}=0$
+It is well known, using a straightening lemma, that if $f,g\in C^{1}$ and $\{f,g\}=0$ in a neighborhood of some $x_{*}\in\mathbb{R}^{2}$, while $\nabla f(x_{*})\not=0,$ then we can write $g=F(f)$ for some $F\in C^{1}$ in a sufficiently small neighborhood of $x_{*}.$ This is a simple application of the inverse function theorem.
 It is also clear that this need not be the case globally.
-There are two important examples one should keep in mind in this regard, both of which we give on $\mathbb{T}^{2}$ (though the examples can be easily modified to work on the unit disk, for example).
+There are two important examples one should keep in mind in this regard, both of which we give in $\mathbb{T}^{2}$ (although the examples can be easily modified to work on the unit disk, for example).
 Example 1: Non-Analytic $g$.
 Take $f(x,y)=\sin(x)\sin(y).$ In this case, each regular level set of $f$ has two connected components consisting of simple closed curves.
-We may take $g$ to simply be a bump function that is constant along the level sets of $f$ restricted to the first quadrant and also compactly supported inside the first quadrant.
+We may take $g$ to be a bump function that is constant along the level sets of $f$ restricted to the first quadrant and also compactly supported inside the first quadrant.
 In this case, $g$ cannot be expressed as a function of $f$ globally on $\mathbb{T}^{2}$ (since $f$ takes the same values in the first and third quadrants, but $g$ takes different values there).
-We give two illustrative figures as follows.
-Figure 1: 2D Contour plot of $g$
-Figure 2: 2D Contour plot of $f$
+We give two illustrative figures as in Figure 1.
+Figure 1. 2D Contour plot of $g$ and $f$ in Example 1
 Example 2: A wall of critical points.
-Take $f=\sin^{2}(x)\sin^{2}(y)$ and $g=\sin(x)\sin(y).$ Then, $g$ cannot be written as $F(f).$ In this case, $f$ and $g$ are both analytic; however, the set of critical points of $f$ partitions the domain into multiple cells surrounded by walls of critical points.
-We give two illustrative figures below.
-The red lines in the figure below are the set of critical points of $f$.
-Figure 3: 2D contour plot of $g$
-Figure 4: 2D contour plot of $f$
+Take $f=\sin^{2}(x)\sin^{2}(y)$ and $g=\sin(x)\sin(y),$ then $g$ cannot be written as $F(f).$ In this case, $f$ and $g$ are both analytic; however, the set of critical points of $f$ partitions the domain into multiple cells surrounded by walls of critical points.
+We give two illustrative figures in Figure 2, where the red lines are the set of critical points of $f$.
+Figure 2. 2D contour plot of $g$ and $f$ in Example 2
 It turns out that these examples encapsulate the only two obstructions to a global functional relation between $f$ and $g.$ We now state a simple consequence of our key proposition in this regard.
 Lemma 1.3.
-Assume that $f,g$ are analytic functions on some open, connected, domain $\Omega.$ Assume that $\{f,g\}=0$ and that the critical points of $f$ of even order are isolated.
-Then, there exists a continuous function $F:\mathbb{R}\rightarrow\mathbb{R}$ so that
+Assume that $f,g$ are analytic functions on some open, connected domain $\Omega.$ Assume that $\{f,g\}=0$ and that the critical points of $f$ of even order are isolated.
+Then, there exists a continuous function $F:\mathbb{R}\rightarrow\mathbb{R}$ such that
 $g=F(f).$
 In particular, in order that there does not exist a global $F$, $f$ must have “walls” of critical points of even order (see Definition 2.1 for the precise meaning of critical point of even order).
-1.4.2 Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0.$
-Given the preceding discussion, the obstruction to having a global relation between $\Delta\psi$ and $\psi$ in a 2d Euler steady state is the existence of a closed curve of critical points of $\psi$ of even order.
-Since $\psi$ is analytic, we can consider the inner-most such curve, $\Gamma$.
-On the region enclosed by $\Gamma$, called $\Omega_{\Gamma}$, $\psi$ will satisfy a semi-linear elliptic equation with non-linearity $F$ and the full gradient of $\psi$ vanishes on $\Gamma$.
+1.4.2. Properties of the set $\{\nabla\psi=0\}$ when $\{\psi,\Delta\psi\}=0$
+Given the preceding discussion, the obstruction to having a global relation between $\Delta\psi$ and $\psi$ in a 2D Euler steady state is the existence of a curve of critical points of $\psi$.
+Since $\psi$ is analytic, we can consider the innermost such curve, $\Gamma$.
+In other words, in the open region enclosed by $\Gamma$ (possibly with the aid of a portion of the boundary $\partial\Omega$), called $\Omega_{\Gamma},$ $\psi$ has only isolated critical points.
+In $\Omega_{\Gamma}$, $\psi$ will satisfy a semilinear elliptic equation with nonlinearity $F$ and the full gradient of $\psi$ vanishes on $\Gamma$.
 Note that the curve $\Gamma$ may have singularities.
-Now there appear three cases:
+In fact, if $\psi$ cannot be written as a solution of (1.4) globally, one of the following three cases must happen:
   1. (i)
-    $\psi$ vanishes higher than quadratically on $\Gamma$,
+    $\psi$ vanishes higher than quadratically on $\Gamma$;
   2. (ii)
-    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ and the associated non-linearity $F$ is not analytic at $c=\psi(\Gamma)$,
+    $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ and the associated nonlinearity $F$ is not analytic at $c=\psi(\Gamma)$;
   3. (iii)
     $\psi$ vanishes only quadratically (somewhere) on $\Gamma$ with $F$ analytic at $c=\psi(\Gamma).$
-In case (iii), it turns out that $F$ may be extended to a larger region containing $\Gamma$ and eventually we return to cases (i) and (ii).
-In both cases (i) and (ii), because $F$ is necessarily non-analytic at $c=\psi(\Gamma)$, it turns out that $\psi-c$ is necessarily single signed in $\Omega_{\Gamma}$–this is a key aspect of the argument.
-In case (i), then we can apply a known result of Brock [5] to deduce that $\psi$ must be radial; a nice feature of Brock’s result [5] is that it does not require the boundary of the domain to be smooth.
+In both cases (i) and (ii), because $F$ is necessarily non-analytic at $c=\psi(\Gamma)$, it turns out that $\psi-c$ is necessarily single-signed in $\Omega_{\Gamma}$–this is a key aspect of the argument.
+In case (i), we can apply a known result of Brock [6] to deduce that $\psi$ must be radial; a nice feature of Brock’s result [6] is that it does not require the boundary of the domain to be smooth.
 The analysis in case (ii) is more difficult and will be discussed in more detail below.
-We give an illustration depicting the potential complexity of the level-set in the figure below [^2] The solid curves on the figure above are the set of critical points of $\psi$ in an elliptical domain.
-The boundary of region A demonstrates situation (i), where all the points are critical points of degree higher than 2.
-The boundary of region B demonstrates the second case (ii), in the case where all of the points are quadratic.
-Regions C and D could represent the third case, where $F$ could be extended from C to D, while D falls into the case (i) or (ii)..
-Figure 5: Illustrations for a potential critical set of $\psi$
-One snag in case (ii) is that we have to show that this curve $\Gamma$ is actually smooth and that $\psi$ is single signed in its interior.
-For general analytic functions, it is well known that the zero set may be quite complicated and may contain multiple singular points of different types [29, 6], even more so in high dimension.
-Note, for example, that the set of critical points in Example 2 above has a singularity at the origin.
-In the figure above, we have to show that case (ii) above actually corresponds to a region like B in the figure (i.e.
-a region with a smooth boundary).
-A key observation here is that, when we have that $\{\psi,\Delta\psi\}=0,$ the degree of vanishing of $\nabla\psi$ on each connected component of $\Gamma$ will necessarily be constant.
-This is established by using the constancy of $\Delta\psi$ on the level sets of $\psi$.
+We give an illustration depicting the potential complexity of the level-set in Figure 3222The solid curves on the figure above are the set of critical points of $\psi$ in an elliptical domain.
+The boundary of the region $A$ demonstrates situation (i), where all points are critical points of degree higher than 2.
+The boundary of the region $B$ demonstrates the second case (ii), in the case where all points are quadratic.
+The regions $C$ and $D$ could represent the third case, where $F$ could be extended from $C$ to $D$, while $D$ falls into case (i) or (ii)..
+Figure 3. Illustrations for a potential critical set of $\psi$
+One snag in case (ii) is that we have to show that this curve $\Gamma$ is actually smooth and that $\psi$ is single-signed in its interior.
+For general analytic functions, it is well known that the zero set may be quite complicated and may contain multiple singular points of different types [34, 7], even more so in high dimensions.
+Note, for example, that the set of critical points in Example 2 in Section 1.4.1 has a singularity at the origin.
+In Figure 3, we have to show that case (ii) above actually corresponds to a region like $B$ in the figure (i.e., a region with a smooth boundary).
+A key observation here is that, when we have $\{\psi,\Delta\psi\}=0,$ the degree of vanishing of $\nabla\psi$ on each connected component of $\Gamma$ will necessarily be constant.
+This is established by using the constancy of $\Delta\psi$ in the levels of $\psi$.
 Once the constancy of the degree is established, we use an elementary fact: singularities in the zero set of an analytic function occur only at points where the degree of vanishing of the function changes.
-The constancy of the degree of vanishing thus implies that $\Gamma$ is smooth.
-1.4.3 Moving Plane Arguments
-From the preceding two steps, we see that if a steady state $\psi$ does not satisfy the global relation $\Delta\psi=F(\psi)$, up to subtracting a constant, we end up with a simply connected domain $\Omega_{\Gamma}$ with analytic boundary where
+The constant degree of vanishing thus implies that $\Gamma$ is smooth.
+In the case (iii), as $F$ is analytic in the boundary $\partial\Omega_{\Gamma}$, we can find an extension of $F$, such that $\psi$ would satisfy $\Delta\psi=F(\psi)$ in a domain containing $\Omega_{\Gamma}$.
+In practice, we do not work with case (iii) directly.
+Instead, based on the discussions on the case (i) and case (ii), we show if $\psi$ is neither radial nor a global solution of (1.4), then $\psi$ would have a nested family of curves of critical points.
+It then leads to a contradiction.
+1.4.3. Moving Plane Arguments
+For case (ii) of the previous subsection, we see that if a steady state $\psi$ does not satisfy the global relation $\Delta\psi=F(\psi)$, up to subtracting a constant, we end up with a simply connected domain $\Omega_{\Gamma}$ with analytic boundary where
 $\Delta\psi=F(\psi)\quad\text{in}\,\,\Omega_{\Gamma},$
@@
 We may also assume that $F(s)\approx\sqrt{s}$ as $s\rightarrow 0.$ As mentioned above, it is not difficult to show that $\psi$ cannot change sign in $\Omega_{\Gamma}$.
-By scaling, we can assume that $0\leq\psi\leq 1$ Now, the overdetermined problem above seems similar to those considered by Serrin in his landmark paper [39].
-Unfortunately, a key difference here is that we do not have the Lipschitz smoothness of the non-linearity $F,$ which is very important for the moving plane method used in [39].
+By scaling, we can assume that $0\leq\psi\leq 1.$ Now, the overdetermined problem above seems similar to those considered by Serrin in his landmark paper [46].
+Unfortunately, a key difference here is that we do not have the Lipschitz smoothness of the nonlinearity $F,$ which is important for the moving plane method used in [46].
 Indeed, as we have already said, near $s=0,$ $F(s)\approx\sqrt{s}$; moreover, $F$ need not be smooth as $s\rightarrow 1$ (the maximum value of $\psi$) either.
-To deal with the singularity as $s\rightarrow 0,$ we use the fact that we have a lower bound on the growth of $\psi$ in the direction normal to the boundary, since the vorticity is non-zero (and constant) on the boundary.
+To deal with the singularity as $s\rightarrow 0,$ we use the fact that we have a lower bound on the growth of $\psi$ in the direction normal to the boundary, since the vorticity is nonzero (and constant) on the boundary.
 Then we invoke a refinement of the Hopf lemma for elliptic problems on small domains under some growth conditions on the coefficients.
@@
 Remarkably, it turns out that the potential singularity as $s\rightarrow 1$ always has the right sign for the maximum principle (see the proof of Lemma 4.8).
-1.5 Organization of the Paper
-In Section 2, we establish some general results about analytic functions satisfying $\{f,g\}=0$ and recall a few results about the zero sets of analytic functions.
+1.5. Organization of the Paper
+In Section 2, we establish some general results on analytic functions satisfying $\{f,g\}=0$ and recall a few results on zero sets of analytic functions.
 The main consequence of this is that, if an analytic Euler solution $\psi$ does not satisfy an equation of the type (1.4), we are ensured the existence of a level set of $\psi$ on which the gradient vanishes.
@@
 This sets us up with an overdetermined elliptic problem in a subdomain of the original domain.
-At the end of that section, we then establish the rigidity of solutions with curves of critical points with higher than quadratic vanishing.
-In Section 4, we establish the rigidity of solutions with a suitably chosen curve of critical points with quadratic vanishing first by deducing certain regularity properties of the function $F$ (defined in a suitable subdomain) and then by applying the moving plane method.
-In Section 4.3 we give a counterexample to Theorem 1.1 in an analytic non-simply connected bounded domain, hence showing the optimality of the statement.
-2 Facts about $\{f,g\}=0$
+At the end of Section 3, we then establish the rigidity of solutions in a domain whose boundary is a curve of critical points with vanishing greater than quadratic.
+In Section 4, we first establish the rigidity of solutions with a suitably chosen curve of critical points with quadratic vanishing, which is based on certain regularity properties of the function $F$ (defined in a suitable subdomain), and the moving plane method.
+Then, we finish the proof of Theorem 1.1 in Section 4.2 via a contradiction argument.
+In Section 4.3, we give a counterexample to Theorem 1.1 in an analytic non-simply connected bounded domain, hence showing the optimality of the statement.
+2. Facts about $\{f,g\}=0$
 The purpose of this section is to establish Proposition 2.1 as well as to recall some basic facts about the zero sets of analytic functions.
-2.1 Key Facts
-We start by giving a notation for the degree of flatness of an analytic function at a point.
+2.1. Key Facts
+We start by giving a notation for the degree of vanishing of an analytic function at a point.
 Definition 2.1.
-For any non-constant analytic $f:\Omega\rightarrow\mathbb{R},$ and $x_{*}\in\Omega,$ we define $d_{f}(x_{*})$ to be the first $k\geq 1$ for which $D^{k}(f)(x_{*})\not=0.$
-We then introduce notation for the set of critical points of $f$ and those with even vanishing:
+For any non-constant analytic function $f:\Omega\rightarrow\mathbb{R},$ and $x_{*}\in\Omega$, $D^{k}f(x_{*})$ denotes the tensor for the $k$-th order derivatives of $f$ at $x_{*}$, and we define $d_{f}(x_{*})$ to be the first $k\geq 1$ for which $D^{k}f(x_{*})\not=0$.
+We then introduce the notations for the set of critical points of $f$ and those with even degree of vanishing:
 $\mathcal{C}(f)=\{x\in\Omega:d_{f}(x)>1\},\qquad\mathcal{C}_{e}(f)=\{x\in\Omega:d_{f}(x)\in 2\mathbb{N}\},$
@@
 Proposition 2.1.
-Assume that $\{f,g\}=0$ in some open connected set $\Omega\subset\mathbb{R}^{2}.$ Suppose that $\Gamma$ is an analytic curve connecting $x,y\in\Omega$ and that $\Gamma\cap\mathcal{C}_{e}(f)=\emptyset.$ Then, if $f(x)=f(y),$ it must be that $g(x)=g(y).$
-To prove this proposition, let us first establish a generalization of the well-known straightening lemma.
+Assume that $\{f,g\}=0$ in some open connected set $\Omega\subset\mathbb{R}^{2}.$ Suppose that $\Gamma_{0}$ is an analytic curve connecting $x,y\in\Omega$ and that $\Gamma_{0}\cap\mathcal{C}_{e}(f)=\emptyset$.
+If $f(x)=f(y)$, then it must hold that $g(x)=g(y).$
+Note that Lemma 1.3 is a direct corollary of Proposition 2.1. To prove this proposition, let us first establish a generalization of the well-known straightening lemma.
 Lemma 2.2.
@@
 Remark 2.2.
-The case $d=1$, which is that $\nabla f(x_{*})\not=0,$ is the standard straightening lemma which is valid only under the assumptions that $f$ and $g$ are $C^{1}$ in a neighborhood of $x_{*}.$
+The case $d=1$, which is that $\nabla f(x_{*})\not=0,$ is the standard straightening lemma, which is valid under the only assumptions that $f$ and $g$ are $C^{1}$ in a neighborhood of $x_{*}.$
 Remark 2.3.
-Strictly speaking, we do not need the generalization to $d>1$ for the results proven later; however, it is conceptually important in order to understand the difference between “extendible” and “unextendible” domains later.
+Strictly speaking, we do not need the generalization to $d>1$ for the results proven later; however, it is conceptually important in order to understand the difference between the “extendable” and “unextendable” domains later.
 The result may also be of independent interest.
-It should be noted that, in this generalization, analyticity plays a key role with the statement not holding true in the smooth class.
-Proof.
+It should be noted that, in this generalization, analyticity plays a key role, with the statement not holding in the smooth class.
+Proof of Lemma 2.2.
 Without loss of generality, by applying suitable translations and a rotation, set $x_{*}=0$, $f(x_{*})=0,$ and $\partial_{1}^{d}f(0)>0.$ Assume that $d$ is odd.
-The point is that we have that $\partial_{1}f(t,0)>0$ for $t\in[-\epsilon,0)\cup(0,\epsilon]$ for some $\epsilon>0$.
+The point is that $\partial_{1}f(t,0)>0$ for $t\in[-\epsilon,0)\cup(0,\epsilon]$ for some $\epsilon>0$.
 In particular, the function $f(t,0)$ is real-analytic and invertible in $(-\epsilon,\epsilon)$.
@@
 In particular, there exists (a unique) $F$ that is analytic in the variable $s^{1/d}$ for which
-$g(t,0)=F(f(t,0)),$
-for every $t\in(-\frac{\epsilon}{2},\frac{\epsilon}{2}).$ Let $\kappa$ be the radius of convergence of the Puiseux series of $F$ at $0$ We now take $\delta<\frac{\epsilon}{2}$ be so small that $f(B_{\delta}(0))\subset(-\kappa,\kappa).$ We will show that
-$g(x)=F(f(x)),$
-for every $x\in B_{\delta}(0).$ Fix any $x\in B_{\delta}(0).$ First, take $x_{1}\in(-\delta,\delta)\setminus\{0\}$.
+$g(t,0)=F(f(t,0)),\,\,\text{for every}\,\,t\in(-\frac{\epsilon}{2},\frac{\epsilon}{2}).$
+Let $\kappa$ be the radius of convergence of the Puiseux series of $F$ at $0.$ We now take $\delta<\frac{\epsilon}{2}$ be so small that $f(B_{\delta}(0))\subset(-\kappa,\kappa).$ We will show that
+$g(x)=F(f(x)),\,\,\text{for every}\,\,x\in B_{\delta}(0).$
+Fix any $x\in B_{\delta}(0).$ First, take $x_{1}\in(-\delta,\delta)\setminus\{0\}$.
 By assumption, $(x_{1},0)$ is not a critical point of $f$.
-It follows that there is a real analytic $G$ so that $g(x)=G(f(x)),$ for every $x$ in a neighborhood of $(x_{1},0).$ In particular, there is an $\epsilon(x_{1})>0$ such that for all $t\in(x_{1}-\epsilon(x_{1}),x_{1}+\epsilon(x_{1})),$ we have that
+It follows that there is a real analytic $G$ so that $g(z)=G(f(z)),$ for every $z$ in a neighborhood of $(x_{1},0).$ In particular, there is an $\epsilon(x_{1})>0$ such that for all $t\in(x_{1}-\epsilon(x_{1}),x_{1}+\epsilon(x_{1})),$ we have
 $G(f(t,0))=F(f(t,0)).$
@@
 Moreover, $F(f(x_{1},t))$ is also analytic except at the possibly discrete set of points where $f(x_{1},t)=0$, where $F(f(x_{1},t))$ has a convergent Puiseux series (note that $f(x_{1},0)\not=0$).
-It follows that $F(f(x_{1},t))$ is actually analytic on $[0,x_{2}]$ and that $g(x_{1},t)=F(f(x_{1},t))$ for every $t\in[0,x_{2}].$ In particular, $g(x_{1},x_{2})=F(f(x_{1},x_{2})).$
+It follows that $F(f(x_{1},t))$ is actually analytic on $0,x_{2}]$ and that $g(x_{1},t)=F(f(x_{1},t))$ for every $t\in[0,x_{2}].$ In particular, $g(x_{1},x_{2})=F(f(x_{1},x_{2})).$
 We are now in a position to prove Proposition 2.1.
 Proof of Proposition 2.1.
-For any point $x\in\Gamma$, by applying Lemma 2.3 to $f$ and $g$, we can find a neighborhood $U_{x}$ and a Puiseux function $F_{x}$ such that $g=F(f)$ in $U_{x}$.
-As the collection of the open set $U_{x}$ forms an open cover for the compact set $\Gamma$, we can choose $U_{x^{1}},\cdots,U_{x^{m}}$, such that $U_{x^{i}}\cap U_{x^{i+1}}\neq\emptyset$ and $\Gamma\subset\cup_{i=1}^{m}U_{x^{i}}$.
-In the case where $f$ is constant in $\Gamma$, as the pair $(g,f)$ satisfies a function relationship in each $U_{x^{i}}$, we have $g$ is constant in $U_{x^{i}}\cap\Gamma$.
-Notice that $\{U_{x^{i}}\}$ is a finite open cover for $\Gamma$, we have $g$ is constant in $\Gamma$.
-In the case where $f$ is not constant in $\Gamma$, we claim $F_{x^{1}}$ can be extended to be a Puiseux function $F$ in the range of $f([\Gamma])$ with $g=F(f)$ in $\Gamma$.
-We first prove that $F_{x^{1}}$ can be extended to a Puiseux function $F$ in the range $f(U_{x^{1}}\cap\Gamma)\cup f(U_{x^{2}}\cap\Gamma)$, with $g=F(f)$ in $\Gamma\cap(U_{x^{1}}\cup U_{x^{2}})$.
-Here $F_{x^{1}}$ and $F_{x^{2}}$ are two Puiseux functions respectively in the range of $f(U_{x^{1}}\cap\Gamma)$ and $f(U_{x^{2}}\cap\Gamma)$, and the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$.
-As $f$ is not constant in $\Gamma$, the analytic assumption of $f$ and $\Gamma$ implies that $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$ contains an open interval in $\mathbb{R}$.
-Now due to the fact that $f(\Gamma\cap U_{x^{1}})$ and $f(\Gamma\cap U_{x^{2}})$ are connected, the condition that $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(U_{x^{1}}\cap U_{x^{2}})$ implies that the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}})\cap f(\Gamma\cap U_{x^{2}})$.
-Hence, we can extend $F_{x^{1}}$ with the aid of $F_{x^{2}}$ to a Puiseux function $F$ in $f(U_{x^{1}})\cup f(U_{x^{1}})$ and $g=F(f)$ in $\Gamma\cap(U_{x}^{1}\cup U_{x}^{2})$.
-We can iterate the above procedure to extend $F_{x^{1}}$ to a Puiseux function $F$ in the range of $f([\Gamma])$ such that $g=F(f)$ in $\Gamma$ and we now finish the proof.
+For any point $\tilde{x}\in\Gamma_{0}$, by applying Lemma 2.2 to $f$ and $g$, there are a neighborhood $U_{\tilde{x}}$ and a Puiseux function $F_{\tilde{x}}$ such that $g=F(f)$ in $U_{\tilde{x}}$.
+As the collection of the open set $U_{\tilde{x}}$ forms an open cover for the compact set $\Gamma_{0}$, we can choose $U_{x^{1}},\cdots,U_{x^{m}}$, such that $U_{x^{i}}\cap U_{x^{i+1}}\neq\emptyset$ and $\Gamma_{0}\subset\cup_{i=1}^{m}U_{x^{i}}$.
+In the case where $f$ is constant in $\Gamma_{0}$, as the pair $(g,f)$ satisfies a functional relationship in each $U_{x^{i}}$, $g$ is constant in $U_{x^{i}}\cap\Gamma_{0}$.
+Notice that $\{U_{x^{i}}\}$ is a finite open cover for $\Gamma_{0}$, therefore, $g$ is constant in $\Gamma_{0}$.
+In the case where $f$ is not constant in $\Gamma_{0}$, we claim that $F_{x^{1}}$ can be extended to be a Puiseux function $F$ in the range of $f(\Gamma_{0})$ with $g=F(f)$ in $\Gamma_{0}$.
+We first prove that $F_{x^{1}}$ can be extended to a Puiseux function $F$ in the range $f(U_{x^{1}}\cap\Gamma_{0})\cup f(U_{x^{2}}\cap\Gamma_{0})$, with $g=F(f)$ in $\Gamma_{0}\cap(U_{x^{1}}\cup U_{x^{2}})$.
+Here $F_{x^{1}}$ and $F_{x^{2}}$ are two Puiseux functions respectively in the range of $f(U_{x^{1}}\cap\Gamma_{0})$ and $f(U_{x^{2}}\cap\Gamma_{0})$, and the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma_{0}\cap U_{x^{1}}\cap U_{x^{2}})$.
+As $f$ is not constant in $\Gamma_{0}$, the analytic assumption of $f$ and $\Gamma_{0}$ implies that $f(\Gamma_{0}\cap U_{x^{1}}\cap U_{x^{2}})$ contains an open interval in $\mathbb{R}$.
+Now due to the fact that $f(\Gamma_{0}\cap U_{x^{1}})$ and $f(\Gamma_{0}\cap U_{x^{2}})$ are connected, the condition that $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(U_{x^{1}}\cap U_{x^{2}})$ implies that the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma_{0}\cap U_{x^{1}})\cap f(\Gamma_{0}\cap U_{x^{2}})$.
+Hence, we can extend $F_{x^{1}}$ with the aid of $F_{x^{2}}$ to a Puiseux function $F$ in $f(U_{x^{1}})\cup f(U_{x^{1}})$ and $g=F(f)$ in $\Gamma_{0}\cap(U_{x}^{1}\cup U_{x}^{2})$.
+Iterate the above procedure to extend $F_{x^{1}}$ to a Puiseux function $F$ in the range of $f(\Gamma_{0})$ such that $g=F(f)$ in $\Gamma_{0}$.
+Hence the proof of the proposition is completed.
 ∎
-2.2 Zero Sets of Analytic Functions
-In this section, we will briefly discuss the structure of zero set of a real analytic function in two variables.
-We firstly recall a classical result concerning the structure of zero set of real analytic function.
+2.2. Zero Sets of Analytic Functions
+In this section, we will briefly discuss the structure of the zero set of a real analytic function in two variables.
+Let us first recall a classical result concerning the structure of the zero set of a real analytic function.
 Proposition 2.3 (See Corollary 2 in [47]).
 Let $f$ be an analytic function near $x^{0}$ and $f(x^{0})=0$, then one of the following three things must happen.
-  1. 1.
+  1. (1)
     $x^{0}$ is an isolated point in the zero set for $f$.
-  2. 2.
+  2. (2)
     $f\equiv 0$.
-  3. 3.
-    The nodal set $\{f=0\}$ near $x^{0}$ is a collection of an even number of analytic curves branching at $x^{0}$.
+  3. (3)
+    The nodal set $\{f=0\}$ near $x^{0}$ is a collection of an even number of analytic curves that branch out at $x^{0}$.
 It is an immediate corollary of Proposition 2.3 that for any given point $x$, which is not isolated in the nodal set for a non-constant analytic function $f$, there are at least two analytic curves $\Gamma_{1}$, $\Gamma_{2}$ branching at $x$ and consisting of nodal points of $f$.
 In most cases, we cannot glue $\Gamma_{1}$ and $\Gamma_{2}$ together to form a single analytic curve.
-For example, in the case $f(x,y)=x^{2}-y^{3}$, the nodal set has a cusp at the origin.
-However, in the case where $\Gamma_{1}$ consists of critical points of the same degree (in particular, if the degree doesn’t jump up at the common point), $\Gamma_{1}\cup\Gamma_{2}$ is an analytic curve.
+For example, in the case $f(x_{1},x_{2})=x_{1}^{2}-x_{2}^{3}$, the nodal set has a cusp at the origin.
+However, in the case where $\Gamma_{1}$ consists of critical points of the same degree (in particular, if the degree does not jump up at the common point), $\Gamma_{1}\cup\Gamma_{2}$ is an analytic curve.
 Lemma 2.4.
-Let $x^{0}$ be a fixed point and we assume there is an analytic curve $\Gamma_{1}:[0,1]\rightarrow\mathbb{R}^{2}$ with $\Gamma_{1}(0)=x^{0}$, and $d_{f}(x)=m$ for all $x\in\Gamma_{1}([0,1])$.
-Then the level set $\{x|f(x)=f(x^{0})\}$ near $x^{0}$ is an analytic curve $\Gamma$.
-Moreover, for any point $x\in\Gamma$, $d_{f}(x)=m$.
+Let $x^{0}$ be a fixed point and we assume that there is an analytic curve $\Gamma_{1}:[0,1]\rightarrow\mathbb{R}^{2}$ with $\Gamma_{1}(0)=x^{0}$, and $d_{f}(x)=m\in[2,+\infty)$ for all $x\in\Gamma_{1}([0,1])$.
+Then the level set $\{x|f(x)=f(x^{0})\}$ near $x^{0}$ is an analytic curve $\Gamma_{0}$.
+Moreover, we have $\Gamma_{1}\subseteq\Gamma_{0}$ and for any point $x\in\Gamma_{0}$, $d_{f}(x)=m$.
 Proof.
-Without loss of generality, by applying suitable rotation, we assume $x^{0}=0$, $f(0)=0$ and $\partial_{1}^{m}f(0)>0$.
-By applying the analytic inverse function theorem to $\partial_{1}^{m-1}f$ at 00, we have the set $\{\partial_{1}^{m-1}f=0\}$ near 00 is an analytic curve $\Gamma$ passing through 00.
-The existence of the analytic curve $\Gamma_{1}$ consisting of critical points of degree $m$ that starts at 00 ensures that for infinitely many points $x^{i}$ accumulate at 00 in $\Gamma$, we have
-$D^{k}f(x^{i})=0,\text{for $1\leq k\leq m-1$ and $i$. }$ (2.1)
-As $\Gamma$ and $f$ are analytic, we have
-$D^{k}f(x)=0,\text{for $x\in\Gamma$, and $1\leq k\leq m-1$.}$ (2.2)
-Then due to the fundamental theorem of calculus, we have $f=0$ in $\Gamma$.
-Now it suffices to find an open set $\mathcal{O}$ near the origin such that $\Gamma\cap\mathcal{O}$ consists of critical point of degree $m$ and $\{f=0\}\cap\mathcal{O}=\Gamma\cap\mathcal{O}$.
-We notice the condition $\partial_{1}^{m}f(0)>0$ ensures that $(1,0)$ is not a tangent vector of $\Gamma$ at 00.
-Now as $f$ and $\Gamma$ are analytic, we can choose $\mathcal{O}:=(-\epsilon,\epsilon)^{2}$ to be a small cube near 00 with the requirement that all the points $x$ in $\mathcal{O}$ satisfy $\partial_{1}^{m}f>0$ and in addition, for any point $x\in\mathcal{O}\setminus\Gamma$, there is a horizontal line $l_{x}$ connecting $x$ and a point $x_{\Gamma}$ in $\Gamma$.
-Now as $\partial_{1}f(x_{\Gamma})=\cdots=\partial_{1}^{m-1}f(x_{\Gamma})=0$ and $\partial_{1}^{m}f>0$ in $l_{x}$, by applying mean value theorem to $f$ in $l_{x}$, we have $f(x)\neq 0$.
-∎
-3 Structure of the set of critical points
-From here and on, we fix a single analytic steady state $\psi,$ satisfying (1.3) on $\bar{\Omega}.$ Towards proving Theorem 1.1, we will also assume that there is no $F$ for which $\psi$ solves (1.4) on all of $\Omega.$ By Proposition 2.2, as well as Proposition 2.6, it must thus be that $\psi$ has curves of critical points.
+Without loss of generality, by applying a suitable rotation, we assume $x^{0}=0$, $f(0)=0$ and $\partial_{1}^{m}f(0)>0$.
+By applying the analytic inverse function theorem to $\partial_{1}^{m-1}f$ at 0, we have the set $\{\partial_{1}^{m-1}f=0\}$ near 0 is an analytic curve $\Gamma_{0}$ passing through 0.
+The existence of the analytic curve $\Gamma_{1}$ consisting of critical points of degree $m$ that starts at 0 ensures that for infinitely many points $\{x^{i}\}_{i=1}^{\infty}\subset\Gamma_{1}$ accumulate at $0\in\Gamma_{0}$, we have
+(2.1) $D^{k}f(x^{i})=0,\,\,\text{for $1\leq k\leq m-1$ and $i\in\mathbb{N}$. }$
+As $\Gamma_{0}$ and $f$ are analytic, it holds that
+(2.2) $D^{k}f(x)=0,\,\,\text{for $x\in\Gamma_{0}$, and $1\leq k\leq m-1$.}$
+Then due to the fundamental theorem of calculus, we have $f=0$ in $\Gamma_{0}$.
+Now that we have identified an analytic curve $\Gamma_{0}$ passing through the origin, which is a level set of $f$ and contains $\Gamma_{1}$, it remains to find an open neighborhood $\mathcal{O}$ of the origin such that
+$\{f=0\}\cap\mathcal{O}=\Gamma_{0}\cap\mathcal{O}.$
+We notice the condition $\partial_{1}^{m}f(0)>0$ ensures that $(1,0)$ is not a tangent vector of $\Gamma_{0}$ at 0.
+Now, since $f$ and $\Gamma_{0}$ are analytic, we can choose $\mathcal{O}:=(-\epsilon,\epsilon)^{2}$ to be a small cube near 0 with the requirement that all points $x$ in $\mathcal{O}$ satisfy $\partial_{1}^{m}f>0$ and, in addition, for any point $x\in\mathcal{O}\setminus\Gamma_{0}$, there is a horizontal line $l_{x}$ connecting $x$ and a point $x_{\Gamma}$ in $\Gamma_{0}$.
+Now, as
+$\partial_{1}f(x_{\Gamma})=\cdots=\partial_{1}^{m-1}f(x_{\Gamma})=0\quad\text{and}\quad\partial_{1}^{m}f>0\,\,\text{in}\,\,l_{x},$
+applying the mean value theorem to $f$ in $l_{x}$, we have $f(x)\neq 0$. ∎
+3. Structure of the Set of Critical Points
+From here and on, we fix a single analytic steady state $\psi,$ satisfying (1.3) on $\bar{\Omega}.$ Towards proving Theorem 1.1, we will also assume that there is no $F$ for which $\psi$ solves (1.4) on all of $\Omega$.
+By Proposition 2.1, as well as Proposition 2.3, it must thus be that $\psi$ has curves of critical points.
 The purpose of this section is to study these in more detail.
@@
 Lemma 3.1.
-Let $\Gamma$ be a curve of points in $\mathcal{C}.$ Suppose that for some $x_{*}\in\Gamma$, $d_{\psi}(x_{*})=2.$ Then, $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $\Gamma.$
+Let $\Gamma_{0}$ be a curve of points in $\mathcal{C}.$ Suppose that for some $x_{*}\in\Gamma_{0}$, $d_{\psi}(x_{*})=2.$ Then, $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $\Gamma_{0}.$
 Proof.
-By Proposition 2.6, for a non-isolated critical point $x$, we have that $D^{2}\psi(x)=0$ is equivalent to $\Delta\psi=0$; moreover, $\Delta\psi$ is constant along any curve.
-This is because, if we take any analytic curve emanating from $x,$ and denote by $\tau$ and $n$ the tangent and normal vectors to the curve at $x,$ we have that $\partial_{\tau n}\psi(x)=\partial_{\tau\tau}\psi(x)=0.$
+We first choose $\delta$ small, such that the Hessian matrix $D^{2}\psi(z)\neq 0$ for all $z\in B_{\delta}(x_{*})$.
+Based on Proposition 2.3, then there exists a $y\in B_{\delta}(x_{*})\cap\Gamma_{0}$ such that $\Gamma_{0}$ near $y$ is an analytic curve.
+Now pick $t$ and $n$ to be the tangent and normal direction of $\Gamma_{0}$ at $y$, respectively.
+The condition that $\Gamma_{0}$ consists of critical points of $\psi$ implies that $\partial_{tt}\psi(y)=\partial_{tn}\psi(y)=0$.
+Now, as $D^{2}\psi(y)\neq 0$, we have $\Delta\psi(y)=\partial_{tt}\psi(y)+\partial_{nn}\psi(y)\neq 0$.
+As $\{\psi,\Delta\psi\}=0$, we have that $\Delta\psi$ is constant along each connected component of $\{\nabla\psi=0\}$.
+In particular, we have $\Delta\psi(x)\neq 0$ and thus $d_{\psi}(x)=2$ for every $x$ in the connected component of $\{\nabla\psi=0\}$ containing $x_{*}$.
+This finishes the proof of the lemma.
+∎
 Using Lemma 2.4, we get the following corollary.
 Corollary 3.2.
-The points for which $d_{\psi}(x)=2$ consist of a collection of isolated points or isolated analytic closed loops in $\Omega$.
+The points in $\Omega$ for which $d_{\psi}(x)=2$ consist of a collection of isolated points or isolated analytic closed loops in $\Omega$.
+Proof.
+Recall that we are assuming that $\psi$ is analytic on $\bar{\Omega}.$ It follows then from Lemma 2.4 that if a curve $\Gamma_{0}$ of critical points of degree $2$ from the interior were to hit $\partial\Omega$, then the degree at the supposed intersection point, $x\in\partial\Omega$, would have to be at least $3$.
+This would contradict the constancy of $\Delta\psi$ along $\Gamma_{0},$ as in Lemma 3.1.
+∎
 The following result, which will be used later, is of a similar flavor.
 Lemma 3.3.
-Let $\Gamma$ be an analytic curve of points of $\psi$ of degree $d\geq 3$. Then, for all $x\in\Gamma,$
+Let $d\geq 3$ and assume that $\Gamma_{0}$ be an analytic curve for which $d_{\psi}(x)=d,$ for all $x\in\Gamma_{0}$. Then, for all $x\in\Gamma_{0},$
 $d_{\Delta\psi}(x)=d-2.$
 Proof.
-Fix $x\in\Gamma$.
-By the rotational invariance of $\Delta,$ we may assume that the tangent vector to $\Gamma$ at $x$ is $(1,0).$ Now, for each $y\in\Gamma$ we have $D^{d-1}\psi(y)=0$, from which we conclude that $D^{d-1}\partial_{1}\psi(x)=0$.
-It thus follows that $D^{d}\psi(x)\not=0$ is equivalent to $\partial_{2}^{d}\psi(x)\not=0.$ This, in turn, is equivalent to $\partial_{2}^{d-2}\Delta\psi(x)\not=0.$
+It is clear that $D^{i}\Delta\psi(x)=0$ for all $x\in\Gamma_{0}$ and $i\leq d-3$.
+Hence it suffices to check $D^{d-2}\Delta\psi(x)\not=0$ for all $x\in\Gamma_{0}$.
+Fix $x^{0}\in\Gamma_{0}$.
+By the rotational invariance of $\Delta,$ we may assume that the tangent vector to $\Gamma_{0}$ at $x^{0}$ is $(1,0).$ Now, for each $y\in\Gamma_{0}$ we have $D^{d-1}\psi(y)=0$.
+It follows that $D^{d-1}\partial_{1}\psi(x^{0})=0$.
+Now $D^{d}\psi(x^{0})\not=0$ implies that $\partial_{2}^{d}\psi(x^{0})\not=0$.
+Hence, one has $D^{d-2}\Delta\psi(x^{0})\not=0$.
+This concludes the proof.
+∎
 Remark 3.1.
@@
 This is shown in the thesis of the second named author [31].
-3.1 The Inner-Most Loop
-Since $\mathcal{C}(\psi)$ contains curves, $\Omega$ must be partitioned by the level set $\{\nabla\psi=0\}$ (this follows from the above discussion and Proposition 2.6).
-In particular, we can find at least one curve of critical points that is either a closed loop in $\Omega$ or starts and ends at $\partial\Omega.$ Let $\Gamma$ be one such curve with the property that the area it encloses contains no curves of critical points.
+3.1. The Innermost Loop
+Since $\mathcal{C}(\psi)$ contains curves, $\Omega$ must be partitioned by the level set $\{\nabla\psi=0\}$ (this follows from the above discussions and Proposition 2.3).
+In particular, we can find at least one curve of critical points that is either a closed loop in $\Omega$ or starts and ends at $\partial\Omega$.
+Note that in the case where the curve starts and ends at $\partial\Omega$, it must touch the boundary at two points (the two points could be the same).
+Together with a boundary portion between these two points, the curve would still enclose a closed simply connected region.
+Now, let $\Gamma$ be a curve with the property that the region it encloses (possibly with the aid of a portion of the boundary) contains no curves of critical points.
 Let us call this enclosed domain $\Omega_{\Gamma},$ which is necessarily simply connected.
-All of the analysis below will be restricted to $\Omega_{\Gamma}.$ Since $\Omega_{\Gamma}$ only contains isolated critical points of $\psi,$ we have that
-$\Delta\psi=F(\psi)\text{ in $\overline{\Omega_{\Gamma}},$ }$ (3.1)
+All of the analysis below will be restricted to $\Omega_{\Gamma}.$ Since the critical points of $\psi,$ contained in $\Omega_{\Gamma}$ must be isolated, we have that
+(3.1) $\Delta\psi=F(\psi)\text{ in $\overline{\Omega_{\Gamma}},$ }$
 for some continuous function $F$. By subtracting a constant, we may also assume that
-$\psi|_{\partial\Omega_{\Gamma}}=0.$
-We will now show that we have, in fact, that $\nabla\psi|_{\partial\Omega_{\Gamma}}=0,$ which is automatic when $\Gamma$ is a loop.
+$\psi=0\quad\text{on}\,\,{\partial\Omega_{\Gamma}}.$
+We will now show actually $\nabla\psi|_{\partial\Omega_{\Gamma}}=0,$ which is automatic when $\Gamma$ is a loop inside $\Omega$.
+Note that if $\Gamma$ is not a loop in (the interior of) $\Omega$, then its endpoints must lie on $\partial\Omega$ (and may coincide).
 Lemma 3.4.
-If $\Gamma$ touches $\partial\Omega,$ then the portion of $\partial\Omega$ between the two endpoints of $\Gamma$ lies in $\mathcal{C}(\psi).$
+If $\Gamma$ touches $\partial\Omega,$ then the portion of $\partial\Omega$ between the two endpoints of $\Gamma$ in $\partial\Omega_{\Gamma}$ lies in $\mathcal{C}(\psi).$
 Proof.
-By Corollary 3.2, if $\Gamma$ touches $\partial\Omega$, then $d_{\psi}(x)\geq 3$ for all $x\in\Gamma.$ Moreover, $\psi$ must then vanish on $\partial\Omega.$ Now, take some $x\in\partial\Omega.$ If $\nabla\psi(x)\not=0,$ then $F$ from (3.1) must be analytic in a neighborhood of $0$ In addition, the condition that $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$ shows $F(0)=0$.
-This, however, contradicts Lemma 3.3, since $F$ being analytic with $F(0)=0$ would imply that $\Delta\psi=F(\psi)$ vanishes to the same order (or higher) than $\psi$ on $\Gamma.$
+First, we claim that if $\Gamma$ touches $\partial\Omega$, then $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$.
+Suppose that there is a point $x_{0}\in\Gamma\cap\Omega$ where $\psi$ vanishes at degree $2$.
+By Lemma 3.1, all the points in $\Gamma$ have to be critical points of $\psi$ with vanishing degree $2$.
+However, just as in the proof of Corollary 3.2, we have shown that $\Gamma$ cannot touch the boundary.
+This leads to a contradiction.
+Second, $\nabla\psi$ must vanish on $\partial\Omega\cap\partial\Omega_{\Gamma}$.
+Indeed, take some $x\in\partial\Omega\cap\partial\Omega_{\Gamma}$, if $\nabla\psi(x)\not=0,$ then it follows from the inverse function theorem that $F$ from (3.1) must be analytic in a neighborhood of 0 (as $\psi$ takes zero value in $\partial\Omega$).
+In addition, the condition that $d_{\psi}(x)\geq 3$ for all $x\in\Gamma$ shows $\psi$ takes the value 0 in $\Gamma$ and that $F(0)=0$.
+Now as $\Delta\psi=F(\psi)$ with $F$ analytic at 0 and $F(0)=0$, considering the leading term of $F(\psi)$ near $\Gamma$, the vanishing degree of $\Delta\psi$ at $\Gamma$ is greater than or equal to the vanishing degree of $\psi$ at $\Gamma$.
+This would contradict Lemma 3.3.
+Hence, the proof of the lemma is complete.
+∎
+Based on Lemmas 3.4 and 3.2, we have the following corollary concerning $\partial\Omega_{\Gamma}$.
 Corollary 3.5.
-One of the following two can hold
-  • Case 1: $\partial\Omega_{\Gamma}$ is analytic and contains only critical points of degree 2.
-  • Case 2: $\partial\Omega_{\Gamma}$ is a closed loop containing critical points of degree larger than 2.
+One of the following two cases must hold.
+  • Case 1: $\partial\Omega_{\Gamma}$ is an analytic loop in $\Omega$ and contains only critical points of degree 2.
+  • Case 2: $\partial\Omega_{\Gamma}$ is a closed loop in $\bar{\Omega}$ containing critical points of degree larger than 2.
 Our goal will now be to show that, in both cases, $\psi$ is necessarily a radial function.
 We will first deal with Case 2, which will be solved essentially by applying existing results from the literature.
-Case 1 will be dealt with in the following section using the moving plane method.
-3.2 Rigidity in Case 2
-In the previous subsection, we reduced the problem into studying an semi-linear elliptic equation of Serrin type [39].
-More precisely, we found a simply connected domain $\Omega_{\Gamma}$ for which $\nabla\psi$ vanishes on $\partial\Omega_{\Gamma}$ and for which $\Omega_{\Gamma}$ contains only isolated critical points of $\psi$.
+Case 1 will be divided into two subcases based on whether the function $F$ appeared in (3.1) is analytic at the value of the stream function on $\partial\Omega_{\Gamma}$.
+These two subcases will be dealt with in the following section using the moving plane method and contradiction argument, respectively.
+3.2. Rigidity in Case 2
+In the previous subsection, we reduced the problem into studying a semilinear elliptic equation of Serrin type [46].
+More precisely, we found a simply connected domain $\Omega_{\Gamma}$ for which $\nabla\psi$ vanishes on $\partial\Omega_{\Gamma}$ and $\Omega_{\Gamma}$ contains only isolated critical points of $\psi$.
 As a consequence of Lemma 1.3, we deduced that $\psi$ solves (3.1) in $\Omega_{\Gamma}$.
-In this section, we consider Case 2 of Corollary 3.6, where all the points of $\partial\Omega_{\Gamma}$ are critical points with degree greater than or equal to $3$.
-Our goal will be to establish the following proposition.
+In this section, we consider Case 2 of Corollary 3.5, where all points in $\partial\Omega_{\Gamma}$ are critical points with degree greater than or equal to $3$.
+Our goal is to establish the following proposition.
 Proposition 3.6.
-Let $\Omega_{\Gamma}$ be a simply connected domain and $\psi$ be an analytic function in $\overline{\Omega_{\Gamma}}$ such that there is no non-isolated critical points of $\psi$ in $\Omega_{\Gamma}\cap\{\psi=0\}$.
-Now we further assume there is a continuous function $F$, such that $\psi$ satisfies the following over-determined elliptic system.
-$\Delta\psi(x)=F(\psi)(x),\text{ for $x\in\Omega_{\Gamma}$,}$ (3.2)
-$\psi(x)=\nabla\psi(x)=\nabla^{2}\psi(x)=0,\text{ for all $x\in\partial\Omega_{\Gamma}$},$
+Let $\Omega_{\Gamma}$ be a simply connected domain and $\psi$ be an analytic function in $\overline{\Omega_{\Gamma}}$ such that there is no non-isolated critical point of $\psi$ in $\Omega_{\Gamma}\cap\{\psi=0\}$.
+Now we further assume that there is a continuous function $F$, such that $\psi$ satisfies the following overdetermined problem for an elliptic equation.
+(3.2) $\Delta\psi(x)=F(\psi)(x),\text{ for $x\in\Omega_{\Gamma}$,}$
+$\psi(x)=\nabla\psi(x)=\nabla^{2}\psi(x)=0,\text{ for all $x\in\partial\Omega_{\Gamma}$}.$
 Then $\psi$ is a radial function.
-We first recall the following result based on Steiner symmetrization method.
+We first recall the following result based on the Steiner symmetrization method.
 Proposition 3.7 (From Theorem 5.2 and Theorem 6.1 in [6] or Theorem 3.2 in [43]).
-Let $\Omega$ be an open bounded set and $\psi$ a non-negative $C^{2}$ solution to
-[…]
-Then $F$ being continuous implies that $\psi$ is locally radial that is:
+Let $\Omega$ be an open bounded domain and $\psi$ be a non-negative $C^{2}$ solution to
+(3.3) $\left\{\begin{aligned} &\Delta{\psi}(x)=F(\psi)(x),\quad&x\in\Omega,\\ &\psi(x)=0,\ \nabla\psi(x)=0,\nabla^{2}\psi(x)=0,\quad&x\in\partial\Omega.\end{aligned}\right$.
+Then $F$ being continuous implies that $\psi$ is locally radial in the sense that
 $\mathbb{D}:=\{0<\psi<sup_{x\in\Omega}\psi(x)\}=A\cup\mathcal{C},$
-where $A=\cup_{i}A_{i}$, $\mathcal{C}=\{\nabla\psi=0\}$ and $A_{k}$ are disjoint open annuli or balls and $\psi$ is radially symmetric in $A_{k}$.
+where $A=\cup_{i}A_{i}$, $\mathcal{C}=\{\nabla\psi=0\}$ and $A_{i}$ are disjoint open annuli or balls and $\psi$ is radially symmetric in $A_{i}$.
 Proof of Proposition 3.6.
-Given Proposition 3.8, all we need to do is show that $\psi$ is single signed on $\Omega_{\Gamma}.$ Towards a contradiction, assume that $\psi$ takes both positive and negative values in $\Omega_{\Gamma}.$ Simply by continuity, it follows that the set of zeros of $\psi$ in $\Omega_{\Gamma}$ is infinite.
-It follows, by our choice of $\Gamma,$ that at least one of those zeros is not a critical point of $\psi$ (since the critical set in $\Omega_{\Gamma}\cap\{\psi=0\}$ is discrete).
+Given Proposition 3.7, all we need to do is show that $\psi$ is single-signed in $\Omega_{\Gamma}.$ Towards a contradiction, assume that $\psi$ takes both positive and negative values in $\Omega_{\Gamma}.$ Simply by continuity, it follows that the set of zeros of $\psi$ in $\Omega_{\Gamma}$ is infinite.
+It follows from our choice of $\Gamma$ that at least one of those zeros is not a critical point of $\psi$ (since the critical set in $\Omega_{\Gamma}\cap\{\psi=0\}$ is discrete).
 It follows that $F$ is analytic in a neighborhood of zero.
-This contradicts Lemma 3.3 (just like in the proof of Lemma 3.5).
+This contradicts Lemma 3.3 (just as in the proof of Lemma 3.4).
 ∎
-4 Moving Plane Argument
-We now move on to prove the rigidity in Case 1 of Corollary 3.6.
-In particular, we may assume that there exists a simply connected domain $\Omega_{\Gamma}$ with $\partial\Omega_{\Gamma}$ analytic with $\psi$ satisfying
-$\Delta\psi(x)=F(\psi)(x),\text{for $x\in\Omega_{\Gamma}$,}$ (4.1)
+4. Moving Plane Arguments and the Proof of the Main Theorem
+We now move on to prove the rigidity in Case 1 of Corollary 3.5, where our basic assumption is that there is no $F$ for which $\psi$ solves (1.4) in all of $\Omega$.
+More precisely, we may assume that there exists a simply connected domain $\Omega_{\Gamma}$ with $\partial\Omega_{\Gamma}$ analytic and $\psi$ satisfies
+(4.1) $\Delta\psi(x)=F(\psi)(x),\text{for $x\in\Omega_{\Gamma}$,}$
 $\psi(x)=\nabla\psi(x)=0,\text{for all $x\in\partial\Omega_{\Gamma}$}.$
 Moreover, by scaling, we may assume that $F(0)=1.$ In particular,
-$\partial_{nn}\psi|_{\partial\Omega_{\Gamma}}=1.$
-An important consequence is that $\psi$ is increasing in the direction normal to the boundary, in a tubular neighborhood of the boundary.
+$\partial_{nn}\psi=1\quad\text{on}\,\,\partial\Omega_{\Gamma}.$
+An important consequence is that $\psi$ increases in the normal direction of the boundary, in a tubular neighborhood of the boundary.
 In the case where $F$ is $C^{1}$ and $\psi$ is non-negative, we have the following rigidity result due to Serrin [46] (see also the nice review [41]).
-Theorem 4.1 (Serrin, 1971).
-Assume that $\psi$ satisfies (4.1) and that $\psi>0$ in $\Omega_{\Gamma},$ while $F$ is Lipschitz continuous. Then, $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
+Theorem 4.1 (Serrin, [46]).
+Assume that $\psi$ satisfies (4.1) and that $\psi>0$ in $\Omega_{\Gamma},$ while $F$ is Lipschitz continuous. Then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
 In our case, we cannot assume that $F$ is $C^{1}$ in general. In fact, there are two cases.
 Lemma 4.2.
-It is either that
-  • $F(s)$ is not analytic at $s=0,$ in which case $\psi$ is non-negative on $\Omega_{\Gamma}.$ This is called the unextendable case.
+One of the following two cases must happen.
+  • $F(s)$ is not analytic at $s=0,$ in which case $\psi$ is non-negative in $\Omega_{\Gamma}.$ This is called the unextendable case.
   • $F(s)$ is analytic at $s=0.$ This is called the extendable case.
-4.1 The Unextendable Case
-This is the case where $F(s)$ is not analytic in $s$ at 00.
+Proof.
+To prove the lemma, it suffices to check that $\psi$ is non-negative on $\Omega_{\Gamma}$ when $F$ is not analytic near 0.
+In fact, the proof is quite similar to that in the proof of Proposition 3.6.
+Now we assume, towards a contradiction, that there is a point $x_{*}\in\Omega_{\Gamma}$ with $\psi(x_{*})<0$.
+Since $\nabla\psi=0$ and $\partial_{nn}\psi=1$ on $\partial\Omega_{\Gamma}$, $\psi$ is positive in a tubular neighborhood of $\Gamma$ excluding $\Gamma$ itself.
+Hence, by continuity, it follows from $\psi(x_{*})<0$ that there are infinitely many zeros of $\psi$ in $\Omega_{\Gamma}$.
+Since $F$ is not analytic at $s=0,$ the zeros of $\psi$ in $\Omega_{\Gamma}$ must be critical points.
+In particular, $\psi$ must have infinitely many critical points in $\Omega_{\Gamma}$.
+This contradicts the choice of $\Gamma$ as the innermost loop.
+∎
+4.1. The Unextendable Case
+This is the case where $F(s)$ is not analytic in $s$ at 0.
 As mentioned in Lemma 4.2, $\psi$ must be non-negative in $\Omega_{\Gamma}$, since otherwise there would be infinitely many points where $\psi=0$ in $\Omega_{\Gamma}$.
@@
 Analogously as in the proof of Proposition 3.6, the condition that $\psi$ does not have a non-isolated critical point in $\Omega_{\Gamma}$ implies that $F$ is analytic in $(0,1)$.
-Our goal in this section is to prove the following proposition:
+Our goal in this subsection is to prove the following proposition.
 Proposition 4.3.
-Let $\Omega_{\Gamma}$ be a simply connected analytic domain, $\psi$ be a non-negative analytic function in $\overline{\Omega_{\Gamma}}$.
-Assume $F\in C([0,1])$ is analytic on $(0,1).$ If $F$ is not analytic at 00 and $\psi$ satisfies the over-determined elliptic equation (4.1), then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
+Let $\Omega_{\Gamma}$ be a simply connected analytic domain, and $\psi$ be a non-negative analytic function in $\overline{\Omega_{\Gamma}}$.
+Assume $F\in C([0,1])$ is analytic on $(0,1).$ If $F$ is not analytic at 0 and $\psi$ satisfies the over-determined problem (4.1) for elliptic equation, then $\Omega_{\Gamma}$ is a ball and $\psi$ is radial.
 Remark 4.1.
-Without a loss of generality, we may assume $F(0)=1,$ since the case $F(0)=0$ is already covered by Proposition 3.7.
+Without loss of generality, we may assume $F(0)=1,$ since the case $F(0)=0$ has already been covered by Proposition 3.6.
 For the proof of Proposition 4.3, it suffices to prove that $\psi$ is even symmetric to all directions $e\in\mathbb{S}^{1}$.
-Without loss of generality, by rotation, $e=(1,0)$.
-We will use the moving plane method to show that $\psi$ is even symmetric to $(1,0)$.
+Without loss of generality, by rotation, we need only to prove the symmetry with respect to $e=(1,0)$.
+The key is to use the moving plane method to show that $\psi$ is even symmetric to $(1,0)$.
 We start with a series of lemmas concerning the regularity of $F$ and the critical set of $\psi$.
@@
 Lemma 4.4.
-When the maxima of $\psi$ are critical points of degree $2k_{0},$ for some $k_{0}\geq 1,$ $F(s)$ admits the following convergent expansion.
-$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{ for $0\leq 1-s\ll 1$,with $a_{k}$ real and $a_{k_{0}-1}<0$. }$ (4.2)
+When the maxima of $\psi$ are critical points of degree $2k_{0},$ for some $k_{0}\geq 1,$ $F(s)$ admits the following convergent expansion, for $0\leq 1-s\ll 1$,
+(4.2) $F(s)=\sum_{k=k_{0}-1}^{\infty}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{ with $a_{k}$ real and $a_{k_{0}-1}<0$. }$
 Remark 4.2.
@@
 In the case $d_{\psi}(x^{0})=2$, without loss of generality, by a suitable rotation and translation, we assume $x^{0}=0$ and $\partial_{11}\psi(0)<0$.
-As $\partial_{1}\psi(0)=0$ and $\partial_{11}\psi(0)<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$ and $f_{-}(s)$, with $f_{+}(s)\geq 0$ and $f_{-}(s)\leq 0$.
-We show that $f_{+}$ and $f_{-}$ admit a Puisuex expression near $1$.
+As $\partial_{1}\psi(0)=0$ and $\partial_{11}\psi(0)<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 0 has exactly two inverses $f_{+}(s)$ and $f_{-}(s)$, with $f_{+}(s)\geq 0$ and $f_{-}(s)\leq 0$.
+We show that $f_{+}$ and $f_{-}$ admit a Puisuex expansion near $1$.
 In particular, we will show that there is an analytic function $G$, such that $f_{+}(s)=G(\sqrt{1-s})$ and $f_{-}(s)=G(-\sqrt{1-s})$.
-Indeed given the Taylor expression of $\psi(t,0)=1+a_{2}t^{2}+\sum_{n\geq 3}a_{n}t^{n}$ with $a_{2}<0$, we have
-$\psi(f_{\pm},0)=s\Leftrightarrow f_{\pm}^{2}\left(1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}\right)=\frac{s-1}{a_{2}}\Leftrightarrow f_{\pm}\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}f_{\pm}^{k-2}}=\pm\sqrt{\frac{s-1}{a_{2}}}.$ (4.3)
-Now the function $s\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}s^{k-2}}$ is an analytic function with a non-vanishing derivative at 00, hence it admits an analytic inverse $H$ and we can correspondingly choose $G(s)=H(\frac{s}{\sqrt{-a_{2}}})$.
-With the aid of $f_{+}$, $f_{-}$ and $G$ in the claim, for $0\leq 1-s\ll 1$, we have
+Indeed, given the Taylor expression of $\psi(t,0)=1+a_{2}t^{2}+\sum_{n\geq 3}a_{n}t^{n}$ with $a_{2}<0$, we have
+(4.3) $\psi(f_{\pm}(s),0)=s\Leftrightarrow$ $f_{\pm}^{2}(s)\left(1+\sum_{k\geq 3}\frac{a_{k}}{a_{2}}f_{\pm}^{k-2}(s)\right)=\frac{s-1}{a_{2}}$
+$\Leftrightarrow$ $f_{\pm}(s)\sqrt{1+\sum_{k\geq 3}\frac{a_{k}}{a_{2}}f_{\pm}^{k-2}(s)}=\pm\sqrt{\frac{s-1}{a_{2}}}.$
+Note that the function $s\sqrt{1+\sum_{k\geq 3}\frac{a_{3}}{a_{2}}s^{k-2}}$ is an analytic function with a non-vanishing derivative at 0, hence it admits an analytic inverse $H$.
+Therefore, we can correspondingly choose $G(s)=H(\frac{s}{\sqrt{-a_{2}}})$ so that $f_{\pm}(s)=G(\pm\sqrt{1-s})$.
+With the aid of $_{+}, ff_{-}$, and $G$, for $0\leq 1-s\ll 1$, we have
 $F(s)=\Delta\psi(G(\sqrt{1-s}),0)=\Delta\psi(G(-\sqrt{1-s}),0).$
-The fact that $\Delta\psi(G(\sqrt{1-s}),0)$ agrees with $\Delta\psi(G(-\sqrt{1-s}),0)$ shows that $F$ is analytic near $1$.
-Now in the case where $d_{\psi}(x^{0})=2k_{0}\geq 4$, again assume that $x^{0}=0$ and $\partial_{1}^{2k_{0}}\psi<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 00 has exactly two inverses $f_{+}(s)$, $f_{-}(s)$.
-Similarly as in the case where $d_{\psi}(x^{0})=2$, we can show that $F$ near $1$ has the Puiseux expression.
-$F(s)=\sum_{k=k_{0}-1}a_{k}(1-s)^{\frac{k}{k_{0}}},\text{with $a_{k}$ real, $a_{k_{0}-1}<0$, for $0\leq 1-s\ll 1$. }$ (4.4)
-Analogously, by taking a normal line segment through a boundary point $x\in\partial\Omega_{\Gamma}$, we have $F(s)$ is analytic in the variable $\sqrt{s}$ near 00.
-Together with the condition that $F$ is not analytic near 00, we have the following Puiseux expression of $F$ near 00.
+The fact that $\Delta\psi(G(\sqrt{1-s}),0)$ agrees with $\Delta\psi(G(-\sqrt{1-s}),0)$, implies that all the terms consisting of odd powers of $\sqrt{1-s}$ vanish in the Puiseux expansion of $F$ near $1$.
+Therefore, $F$ is analytic near $1$.
+Now in the case where $d_{\psi}(x^{0})=2k_{0}\geq 4$, again assume that $x^{0}=0$ and $\partial_{1}^{2k_{0}}\psi<0$, for $0\leq 1-s\ll 1$, the function $\psi(t,0)$ near 0 has exactly two inverses $f_{+}(s)$, $f_{-}(s)$.
+Similarly, as in the case where $d_{\psi}(x^{0})=2$, it follows that $F$ near $1$ has the Puiseux expansion.
+More precisely, for $0\leq 1-s\ll 1$, we have (4.2).
+This finishes the proof of the lemma.
+∎
+Analogously, by taking a normal line segment through a boundary point $x\in\partial\Omega_{\Gamma}$, we have $F(s)$ is analytic in the variable $\sqrt{s}$ near 0.
+Together with the condition that $F$ is not analytic near 0, we have the following Puiseux expansion of $F$ near 0.
 Lemma 4.5.
-$F$ near 00 admits a convergent Puiseux series expression:
-$F(s)=\sum_{k=0}a_{k}s^{\frac{k}{2}},\text{for $0\leq s\ll 1$ , with $a_{k}$ real, $a_{0}=1$ and $a_{k_{0}}\neq 0$ for an odd number $k_{0}$.}$ (4.5)
-Moreover, by arguing similarly as in the case $d_{\psi}(x^{0})=2$ in the proof of Lemma 4.5, the condition $F$ is not analytic at 00 implies that $\psi$ is positive in $\Omega_{\Gamma}$.
-Together with fact
+$F$ near 0 admits a convergent Puiseux series expression: for $0\leq s\ll 1$, one has
+(4.4) $F(s)=\sum_{k=0}^{\infty}a_{k}s^{\frac{k}{2}},$
+where $a_{k}\in\mathbb{R}$, $a_{0}=1$ and $a_{k_{0}}\neq 0$ for an odd number $k_{0}$.
+Moreover, arguing similarly as in the case $d_{\psi}(x^{0})=2$ in the proof of Lemma 4.4, the condition $F$ is not analytic at 0 implies that $\psi$ is positive in $\Omega_{\Gamma}$.
+Together with the fact
 $\partial_{nn}\psi|_{\partial\Omega_{\Gamma}}=1,$
-we can show $\psi$ has a uniform lower bound in $\Omega_{\Gamma}$.
+we can show that $\psi$ has a uniform lower bound in $\Omega_{\Gamma}$.
 Lemma 4.6.
-There exists $C>0$ such that for all $x\in\Omega_{\Gamma}$,
-$\frac{dist(x,\partial\Omega_{\Gamma})^{2}}{C}\leq\psi(x)\leq Cdist(x,\partial\Omega_{\Gamma})^{2}.$ (4.6)
+There exists a constant $C>0$ such that for all $x\in\Omega_{\Gamma}$,
+(4.5) $\frac{dist(x,\partial\Omega_{\Gamma})^{2}}{C}\leq\psi(x)\leq Cdist(x,\partial\Omega_{\Gamma})^{2}.$
 Now we are ready to apply the moving plane method to show that $\psi$ is even symmetric in the direction $e=(1,0)$.
 Without loss of generality, by translation, we fix $\inf_{x\in\Omega_{\Gamma}}x_{1}=0$.
-For $\lambda>0$, let $H_{\lambda}:=\{x_{1}=\lambda\}$.
-We now reflect $\psi$ and $\Omega_{\Gamma}$ through the hyperplane $H_{\lambda}$:
-$\pi_{\lambda}(x_{1},x_{2}):=(2\lambda-x_{1},x_{2}),\ \psi_{\lambda}(x):=\psi(\pi_{\lambda}(x)),\ \Omega_{\lambda}:=\{x_{1}<\lambda\}\cap\Omega_{\Gamma},\text{ and }\breve{\Omega}_{\lambda}:=\pi_{\lambda}(\Omega_{\lambda}).$
+For $\lambda>0$, define
+$H_{\lambda}:=\{x_{1}=\lambda\},\,\,\Omega_{\lambda}:=\{x_{1}<\lambda\}\cap\Omega_{\Gamma},\text{ and }\pi_{\lambda}(x_{1},x_{2}):=(2\lambda-x_{1},x_{2}).$
+We now reflect $\psi$ and $\Omega_{\lambda}$ through the hyperplane $H_{\lambda}$ as follows:
+$\psi_{\lambda}(x):=\psi(\pi_{\lambda}(x)),\,\,\breve{\Omega}_{\lambda}:=\pi_{\lambda}(\Omega_{\lambda}).$
 We are going to show $\psi_{\lambda}$ is always less than or equal to $\psi$ in $\breve{\Omega}_{\lambda}$ before $\partial\breve{\Omega}_{\lambda}$ touches the physical boundary $\partial\Omega_{\Gamma}$ tangentially.
@@
 Since $\partial\Omega_{\Gamma}$ is analytic, for $\lambda\ll 1$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$.
-Moreover, as $\nabla\psi=0$ and $\partial_{nn}\psi=1$ in $\Omega_{\Gamma}$, $\psi$ satisfies the reflection monotonicity property at the level $0<\lambda\ll 1$ We will now prove that the reflection monotonicity property persists unless $\partial\breve{\Omega}_{\lambda}$ hits $\partial\Omega_{\Gamma}$ tangentially.
+Moreover, as $\nabla\psi=0$ and $\partial_{nn}\psi=1$ in $\partial\Omega_{\Gamma}$, $\psi$ satisfies the reflection monotonicity property at the level $0<\lambda\ll 1$ .
+We will now prove that the reflection monotonicity property persists unless $\partial\breve{\Omega}_{\lambda}$ hits $\partial\Omega_{\Gamma}$ tangentially.
 Lemma 4.7.
 Define
-$\lambda_{0}=\sup\{\lambda|\text{$\psi$ satisfies the reflection monotonicity property at level $\lambda$}\},$
+[…]
 We have $h_{\lambda_{0}}=\psi-\psi_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$. Moreover, at least one of the following occurs when $\lambda=\lambda_{0}$.
@@
 In particular, the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is $(0,1)$.
-Figure 6: Internal tangency
-Figure 7: Boundary tangency
-The proof of Lemma 4.10 is based on the observation that as long as $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$, the difference $\psi-\psi_{\lambda}$ satisfies a singular elliptic equation where the singular coefficent has a uniform upper bound.
+Figure 4. Internal tangency and boundary tangency
+The proof of Lemma 4.7 is based on the observation that as long as $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$, the difference $\psi-\psi_{\lambda}$ satisfies an elliptic equation where the coefficient has a uniform estimate.
 Lemma 4.8.
-For all $\lambda>0$, such that $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma},$ $h_{\lambda}:=\psi-\psi_{\lambda}$ satisfies a singular elliptic equation
-$\Delta h_{\lambda}=c_{\lambda}(x)h_{\lambda},$ (4.7)
-where the singular coefficient $c(x)$ has a uniform upper bound
-$c_{\lambda}^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}.$ (4.8)
+For all $\lambda>0$, such that $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma},$ $h_{\lambda}:=\psi-\psi_{\lambda}$ satisfies an elliptic equation
+(4.6) $\Delta h_{\lambda}=c_{\lambda}(x)h_{\lambda},$
+where the coefficient $c(x)$ satisfies the following uniform estimate
+(4.7) $c_{\lambda}^{-}(x):=\max\{-c_{\lambda}(x),0\}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}.$
 Proof.
 Due to the reflection symmetry of the semilinear elliptic equation, $\Delta\psi_{\lambda}=F(\psi_{\lambda})$, and thus
-$\Delta h_{\lambda}=\frac{F(\psi)-F(\psi_{\lambda})}{\psi-\psi_{\lambda}}h_{\lambda}:=c_{\lambda}(x)h_{\lambda}.$ (4.9)
-In the case where $F$ is analytic at $1$, as $F$ is already analytic in $(0,1)$, from Lemma 4.7, there exists a function $G\in C^{1}([0,1])$ and a constant $d$ such that $F(s)=d\sqrt{s}+G(s)$ ($d$ might be zero, which corresponds to the case where the Puiseux expression of $F$ at 00 has no $\sqrt{s}$ term).
+(4.8) $\Delta h_{\lambda}=\frac{F(\psi)-F(\psi_{\lambda})}{\psi-\psi_{\lambda}}h_{\lambda}:=c_{\lambda}(x)h_{\lambda}.$
+In the case where $F$ is analytic at $1$, as $F$ is already analytic in $(0,1)$, from Lemma 4.5, there exists a function $G\in C^{1}([0,1])$ and a constant $d$ such that $F(s)=d\sqrt{s}+G(s)$ ($d$ might be zero, which corresponds to the case where the Puiseux expansion of $F$ at 0 has no $\sqrt{s}$ term).
 Now we have
-$c_{\lambda}=\frac{d}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}$ (4.10)
-Due to the lower bound of $\psi$ in the equation (4.6), we have
-$\psi_{\lambda}>\frac{dist(x,\partial\breve{\Omega}_{\lambda})^{2}}{C},$ (4.11)
-together with the fact that $\psi$ and $\psi_{\lambda}$ enjoy a uniform $C^{1}$ upper bound, we have $|c_{\lambda}|\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, where $M>0$ is independent of $\lambda$.
-In the case where $F$ is not analytic near $1$, since $F$ is analytic in $(0,1)$, Lemma 4.5 and Lemma 4.7 show that there exists an integer $k_{0}\geq 2$, a negative constant $d_{1}<0$, a constant $d_{2}$ and a function $G\in C^{1}([0,1])$, such that $F(s)=d_{1}(1-s)^{\frac{k_{0}-1}{k_{0}}}+d_{2}\sqrt{s}+G(s)$.
+(4.9) $c_{\lambda}=\frac{d}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}$
+It follows from the lower bound of $\psi$ in the estimate (4.5) that one has
+(4.10) $\psi_{\lambda}\geq\frac{dist(x,\partial\breve{\Omega}_{\lambda})^{2}}{C}.$
+This, together with the fact that $\psi$ and $\psi_{\lambda}$ enjoy a uniform $C^{1}$ upper bound, yields $|c_{\lambda}|\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, where $M>0$ is independent of $\lambda$.
+In the case where $F$ is not analytic near $1$, since $F$ is analytic in $(0,1)$, Lemmas 4.4 and 4.5 show that there exist an integer $k_{0}\geq 2$, a negative constant $d_{1}<0$, a constant $d_{2}$, and a function $G\in C^{1}([0,1])$, such that
+$F(s)=d_{1}(1-s)^{\frac{k_{0}-1}{k_{0}}}+d_{2}\sqrt{s}+G(s).$
 We now have the corresponding expression for $c_{\lambda}$
-$c_{\lambda}=d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}+\frac{d_{2}}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}.$ (4.12)
+(4.11) $c_{\lambda}=d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}+\frac{d_{2}}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}.$
 As $d_{1}<0$, the term $d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}$ is non-negative.
-Hence, similiarly as in the case where $F$ is analytic at $1$, we have $c_{\lambda}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, for a positive $M$ independent of $\lambda$.
+Hence, similar to the case where $F$ is analytic at $1$, we have $c_{\lambda}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, for a positive $M$ independent of $\lambda$.
 ∎
-We now prove Lemma 4.10, arguing by contradiction.
+We now prove Lemma 4.7 by the contradiction argument.
 Proof of Lemma 4.7.
@@
 We will show that there is a $\delta\in(0,\delta^{*})$ such that $\psi$ satisfies the reflection monotonicity at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$.
-We start by proving that $h_{\lambda_{0}}$ is positive in $\breve{\Omega}_{\lambda_{0}}$.
-By Lemma 4.11, $h_{\lambda_{0}}$ is a nonnegative solution to $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ with $c^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$.
-Now suppose that there is a point $x^{0}\in\breve{\Omega}$ such that $h_{\lambda}(x^{0})=0$, we automatically have $\nabla h_{\lambda}(x^{0})=0$.
-Then the Proposition A.1 in the appendix shows in a small ball whose boundary contains $x^{0}$, $h_{\lambda_{0}}$ is identically zero.
+We start by proving that $h_{\lambda_{0}}$ is positive in $\breve{\Omega}_{\lambda_{0}}$. By Lemma 4.8, $h_{\lambda_{0}}$ is a nonnegative solution to
+(4.12) $\Delta h_{\lambda_{0}}=c_{\lambda_{0}}(x)h_{\lambda_{0}}\,\,\text{with}\,\,c_{\lambda_{0}}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}.$
+Now suppose that there is a point $x^{0}\in\breve{\Omega}_{\lambda_{0}}$ such that $h_{\lambda_{0}}(x^{0})=0$, we automatically have $\nabla h_{\lambda_{0}}(x^{0})=0$.
+Then the Proposition A.1 in the appendix shows that in a small ball whose boundary contains $x^{0}$, $h_{\lambda_{0}}$ is identically zero.
 Due to the analytic regularity of $\psi$, it further implies that $\psi_{\lambda_{0}}=\psi$.
-Hence, $\Omega_{\Gamma}$ is even symmetric to $H_{\lambda_{0}}$ and we get a contradiction.
-Hence, we now have
-$h_{\lambda_{0}}(x)>0,\text{ for all $x\in\breve{\Omega}_{\lambda_{0}}$.}$ (4.13)
-Anaglously, for any point $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$, we may apply Proposition A.1 to a small ball that hits $x$ boundary-tangentially at $x$, and conclude that
-$\partial_{1}h_{\lambda_{0}}(x)>0,\text{ for all $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$},$ (4.14)
+Hence $\Omega_{\Gamma}$ is even symmetric with respect to $H_{\lambda_{0}}$ and we get a contradiction.
+Therefore, it holds that
+(4.13) $h_{\lambda_{0}}(x)>0,\text{ for all $x\in\breve{\Omega}_{\lambda_{0}}$.}$
+Analogously, for any point $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$, we may apply Proposition A.1 to a small ball that hits the boundary of $\partial\breve{\Omega}_{\lambda_{0}}$ tangentially at $x$, and conclude that
+(4.14) $\partial_{1}h_{\lambda_{0}}(x)>0,\text{ for all $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$},$
 unless $h_{\lambda_{0}}\equiv 0.$
-In the rest of the proof, we will show that there is a $\delta>0$ such that $\psi$ satisfies the reflection monotonicity property at level $\lambda\in[\lambda_{0},\lambda+\delta].$ First, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit $\partial\Omega_{\Gamma}$ boundary-tangentially, for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta^{*}]$, $H_{\lambda}\cap\partial\Omega_{\Gamma}$ consists of two points $x_{\lambda}^{1}$, $x_{\lambda}^{2}$($x_{\lambda}^{1}$ is above $x_{\lambda}^{2}$).
-Moreover, there is a $\epsilon_{0}>0$ such that $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}}^{1})<-\epsilon_{0}$ and $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}})(x_{\lambda_{0}}^{2})>\epsilon_{0}$.
-As $\partial_{12}h_{\lambda}(x_{\lambda}^{1})$ and $\partial_{12}h_{\lambda}(x_{\lambda}^{2})$ are uniformly continuous in $\lambda$ , we choose $\delta_{1}>0$ such that
-$\partial_{12}h_{\lambda}(x_{\lambda}^{1})<-\frac{1}{2}\epsilon_{0},\partial_{12}h_{\lambda}(x_{\lambda}^{2})>\frac{1}{2}\epsilon_{0},\text{ for$\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta_{1}$}.$ (4.15)
+In the rest of the proof, we will show that there is a $\delta>0$ such that $\psi$ satisfies the reflection monotonicity property at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta].$ First, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit the boundary $\partial\Omega_{\Gamma}$ tangentially, for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta^{*}]$, $H_{\lambda}\cap\partial\Omega_{\Gamma}$ consists of two points $x_{\lambda}^{1}$, $x_{\lambda}^{2}$($x_{\lambda}^{1}$ is above $x_{\lambda}^{2}$).
+Moreover, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit the boundary $\partial\Omega_{\Gamma}$ tangentially at the points $x_{\lambda}^{1}$ and $x_{\lambda}^{2}$, the condition that $\partial_{nn}\psi=1,\psi_{tt}\psi=\psi_{tn}=0$ ($t,n$ here represent the tangent and normal direction, respectively) for all points in $\partial\Omega_{\Gamma}$ implies that there is an $\epsilon_{0}>0$ such that
+$\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}}^{1})<-\epsilon_{0}\,\,\text{and}\,\,\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}})(x_{\lambda_{0}}^{2})>\epsilon_{0}.$
+As $\partial_{12}h_{\lambda}(x_{\lambda}^{1})$ and $\partial_{12}h_{\lambda}(x_{\lambda}^{2})$ are uniformly continuous in $\lambda$, there exist a $\delta_{1}>0$ such that
+(4.15) $\partial_{12}h_{\lambda}(x_{\lambda}^{1})<-\frac{1}{2}\epsilon_{0}\,\,\text{and}\,\,\partial_{12}h_{\lambda}(x_{\lambda}^{2})>\frac{1}{2}\epsilon_{0},\text{ for $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta_{1}$}.$
 As $x_{\lambda}^{1}$ is a critical point of $\psi$ and $\psi_{\lambda}$, we have
-$\nabla h_{\lambda}(x_{\lambda}^{1})=0,$ (4.16)
-moreover by symmetry, we have
-$\partial_{11}h_{\lambda}(x_{\lambda}^{1})=\partial_{22}h_{\lambda}(x_{\lambda}^{1})=0.$ (4.17)
-Due to the smoothness of $\partial\Omega_{\Gamma}$ and that the boundary tangency doesn’t occur when $\lambda=\lambda_{0}$, there exists a $\epsilon>0$, such that the direction $(1,-\epsilon)$ is always pointing to the exterior of $\breve{\Omega}_{\lambda}$ at $x_{\lambda}^{1}$ for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$(we may decrease the value of $\delta_{1}$ when it is necessary).
-Now since $D^{3}h_{\lambda}$ is uniformly bounded, from the equations (4.15), (4.16) and (4.17), there exists $r_{1}>0$, such that $h_{\lambda}>0$ in $B_{r_{1}}(x_{\lambda}^{1})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$.
-Analogously, in $B_{r_{1}}(x_{\lambda}^{2})$ (we decrease the value of $\delta_{1}$ when necessary), $h_{\lambda}\geq 0$ in $B_{r_{1}}(x_{\lambda}^{2})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$.
-Now due to the equation (4.14), there is a $\epsilon_{1}>0$ such that we have $\partial_{1}h_{\lambda_{0}}(x)>\epsilon_{1}$ for all $x\in H_{\lambda_{0}}\setminus(B_{r_{1}}(x_{\lambda_{0}}^{1})\cup B_{r_{1}}(x_{\lambda_{0}}^{2}))$.
-Similar to the previous case, we can find a $r_{2}>0$ and a $\delta_{2}\in(0,\delta_{1})$, such that $h_{\lambda}\geq 0$ in the set $T_{\lambda}:\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r_{2}\}\setminus(B_{r}(x_{\lambda}^{1})\cup B_{r}(x_{\lambda}^{2}))\}$ for all $\lambda_{0}\leq\lambda\leq\lambda+\delta_{2}$.
-In summary, we show there is a uniform $r>0$ and $\delta>0$, such that if $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$, we have $h_{\lambda}\geq 0$ in $N_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r\}$, which is a tubular neighborhood of $H_{\lambda}$.
-Note by (4.13) and Lemma 4.8, there is a $\epsilon_{2}>0$, such that $h_{\lambda_{0}}>\epsilon_{2}$ on the set $\breve{\Omega}_{\lambda_{0}}\setminus N_{\lambda_{0}}$.
+(4.16) $\nabla h_{\lambda}(x_{\lambda}^{1})=0.$
+Moreover, by symmetry, it holds that
+(4.17) $\partial_{11}h_{\lambda}(x_{\lambda}^{1})=\partial_{22}h_{\lambda}(x_{\lambda}^{1})=0.$
+Due to the smoothness of $\partial\Omega_{\Gamma}$ and that the boundary tangency doesn’t occur when $\lambda=\lambda_{0}$, there exists an $\epsilon>0$, such that the direction $(1,-\epsilon)$ is always pointing to the exterior of $\breve{\Omega}_{\lambda}$ at $x_{\lambda}^{1}$ for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$(we may decrease the value of $\delta_{1}$ when it is necessary).
+Now since $\partial^{3}h_{\lambda}$ is uniformly bounded, from (4.15), (4.16), and (4.17), there exists an $r_{1}>0$ such that $h_{\lambda}>0$ in $B_{r_{1}}(x_{\lambda}^{1})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$.
+Analogously, $h_{\lambda}\geq 0$ in $B_{r_{1}}(x_{\lambda}^{2})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$ (we decrease the value of $\delta_{1}$ when necessary).
+It follows from the estimate (4.14) that there is an $\epsilon_{1}>0$ such that we have $\partial_{1}h_{\lambda_{0}}(x)>\epsilon_{1}$ for all $x\in H_{\lambda_{0}}\setminus(B_{r_{1}}(x_{\lambda_{0}}^{1})\cup B_{r_{1}}(x_{\lambda_{0}}^{2}))$.
+Similar to the previous case, we can find $r_{2}>0$ and $\delta_{2}\in(0,\delta_{1})$, such that for all $\lambda\in(\lambda_{0},\lambda_{0}+\delta_{2})$, $h_{\lambda}\geq 0$ in the set
+$T_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},\,dist(x,H_{\lambda})<r_{2}\}\setminus(B_{r_{1}}(x_{\lambda}^{1})\cup B_{r_{1}}(x_{\lambda}^{2})).$
+In summary, we show there are a uniform $r>0$ and $\delta>0$, such that if $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$, we have $h_{\lambda}\geq 0$ in $N_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r\}$, which is a tubular neighborhood of $H_{\lambda}$.
+Note by (4.13) and Lemma 4.6, there is an $\epsilon_{2}>0$, such that $h_{\lambda_{0}}>\epsilon_{2}$ on the set $\breve{\Omega}_{\lambda_{0}}\setminus N_{\lambda_{0}}$.
 Now by the uniform continuity of $h_{\lambda}$, we may decrease the value of $\delta$ (still make it positive), such that $h_{\lambda}>\frac{\epsilon_{2}}{2}$ on the set $\breve{\Omega}_{\lambda}\setminus N_{\lambda}$ for all $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$.
@@
 Now we apply Hopf’s lemma to prove that $\psi$ and $\Omega_{\Gamma}$ are even symmetric to the $(1,0)$ direction to finish the proof of Proposition 4.3.
-Proof.
-Lemma 4.10 shows there exists $\lambda_{0}>0$ with $h_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$ and at least one of the internal boundary tangency and boundary tangency happens for $\breve{\Omega}_{\lambda_{0}}$.
-In the case where internal tangency occurs at $x^{0}\in\partial\Omega_{\Gamma}$, by Lemma 4.11, $h_{\lambda_{0}}$ is a non-negative solution of the equation $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ in $\breve{\Omega}_{\lambda_{0}}$, with $c^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$.
-Moreover, we have $\nabla h_{\lambda_{0}}(x^{0})=\nabla\psi(x^{0})-\nabla\psi_{\lambda_{0}}(x^{0})=0$.
+Proof of Proposition 4.3.
+Lemma 4.7 shows there exists a $\lambda_{0}>0$ with $h_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$ and at least one of the internal boundary tangency and boundary tangency happens for $\breve{\Omega}_{\lambda_{0}}$.
+In the case where internal tangency occurs at $x^{0}\in\partial\Omega_{\Gamma}$, by Lemma 4.8, $h_{\lambda_{0}}$ is a non-negative solution of the equation (4.12).
+Moreover, it holds that
+$\nabla h_{\lambda_{0}}(x^{0})=\nabla\psi(x^{0})-\nabla\psi_{\lambda_{0}}(x^{0})=0.$
 Now we apply Proposition A.1 for a sufficiently small ball $B\subset\breve{\Omega}_{\lambda_{0}}$ that hits $\breve{\Omega}_{\lambda_{0}}$ internal-tangentially at $x^{0}$, and conclude $h_{\lambda_{0}}=0$.
@@
 Now we consider the case where the boundary of tangency happens at $x^{0}\in H_{\lambda_{0}}\cap\partial\Omega_{\Gamma}$.
-First, similar to above, we have $\nabla h_{\lambda_{0}}(x^{0})=0$.
-In addition, $\partial\breve{\Omega}_{\lambda_{0}}$ forms a right angle at $x^{0}$, and the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is parallel to direction $(0,1)$ In particular, we have $\nabla^{2}h_{\lambda_{0}}(x^{0})=0$.
+First, similar to above, one has $\nabla h_{\lambda_{0}}(x^{0})=0$.
+In addition, $\partial\breve{\Omega}_{\lambda_{0}}$ forms a right angle at $x^{0}$, and the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is parallel to direction $(0,1)$.
+In particular, we have $\nabla^{2}h_{\lambda_{0}}(x^{0})=0$.
 Now analogously as in the case of the internal tangency, by applying Proposition A.2 in the appendix to a small half ball contained in $\breve{\Omega}_{\lambda_{0}}$ that hits $\Omega_{\Gamma}$ boundary-tangentially, we have $h_{\lambda_{0}}=0$.
@@
 ∎
-4.2 The Extendable Case
-Now we assume that $F$ is analytic at $s=0.$ Using the argument in the proof of Proposition 2.2, we see that we can extend the definition of $F$ so that $\psi$ satisfies
-$\Delta\psi=F(\psi)$
-on an extended domain $\tilde{\Omega}_{\Gamma}$, whose boundary is again either an analytic curve of points of degree 2, or a curve of points of degree 3 or higher.
-By the Jordan curve theorem, $\Gamma$ is either in the interior of the $\partial\Omega_{\Gamma}$ or in the exterior.
-In either case, since we are assuming that $F$ cannot be extended to all of $\Omega$, we will hit a curve of critical points along which $F$ is not analytic.
-In that case, we take the smallest curve inscribed in it with that property and we apply either Proposition 3.7 or 4.3 depending on the degree of vanishing of $\nabla\psi$ on the boundary.
-The proof of Theorem 1.1 is now complete.
-4.3 Counterexample
+4.2. The proof of Theorem 1.1
+In this section, we finish the proof of Theorem 1.1.
+We first introduce the following notation.
+Definition 4.4.
+Given a set $U\subset\mathbb{R}^{2}$ and a function $\psi\in C^{2}(U)$, if there exists a continuous function $F$ such that $\Delta\psi=F(\psi)$ in $U$, then $\psi$ is said to be realizable by $F$ in $U$, or we abbreviate as $\psi$ is realizable in $U$.
+If $F$ is extendable in $\Omega_{\Gamma}$ as in Lemma 4.2, since $\Delta\psi-F(\psi)$ is an analytic function in a tubular neighborhood of $\partial\Omega_{\Gamma}$ and $\Delta\psi-F(\psi)=0$ in $\Omega_{\Gamma}$, $\psi$ must be realizable by $F$ in a domain containing $\Omega_{\Gamma}$.
+Furthermore, we have the following proposition.
+Proposition 4.9.
+Assume that $U$ is a simply connected bounded open set, $\psi\in C^{\omega}(\bar{U})$ is constant on $\partial U$}.
+Assume $\psi$ is realizable by $F$ in $U$ and $F$ is not analytic at $c:=\psi(\partial U)$, then $\psi$ is radially symmetric.
+Proof.
+Without loss of generality, we assume $c=0$.
+Since $F$ is not analytic at 0, $\partial U$ consists of critical points of $\psi$.
+By Proposition 2.3, we can find an innermost loop $\Gamma_{0}$ of critical points of $\psi$ where $F$ loses analyticity (i.e., $F$ is analytic at the values of $\psi$ in every curve of critical points contained in the open region enclosed by $\Gamma_{0}$).
+Now $\psi$ must be single-signed in the region $\Omega_{\Gamma_{0}}$ which is enclosed by $\Gamma_{0}$.
+Otherwise, there exist two points $x^{0},x^{1}\in\Omega_{\Gamma_{0}}$ such that $\psi(x^{0})<0$ and $\psi(x^{1})>0$.
+By the continuity of $\psi$, there is a zero point of $\psi$ in every continuous curve that connects $x^{0}$ and $x^{1}$ and is contained in $\Omega_{\Gamma_{0}}$.
+In particular, $\{\psi=0\}$ has infinitely many points in $\Omega_{\Gamma_{0}}$ and there exists a curve $\Gamma_{1}$ consisting of zeros of $\psi$.
+Since $F$ is not analytic at 0, $\Gamma_{1}$ must be a curve of critical points.
+This contradicts the definition of $\Gamma_{0}$.
+Analogously as in Section 3.2 and Section 4.1, applying Proposition 3.6 or Proposition 4.3 depending on the degree of vanishing of $\psi$ in $\Gamma_{0}$, $\psi$ must be a radially symmetric function.
+∎
+To prove Theorem 1.1, we need the following important lemma.
+Lemma 4.10.
+Assume that $U$ is a simply connected bounded open set, $\psi\in C^{\omega}(\bar{U})$ is constant on $\partial U$, and $\{\psi,\Delta\psi\}=0$ in $U$.
+If $\psi$ is neither radially symmetric nor realizable in $U$, then there exists a simply connected domain $\tilde{U}\varsubsetneq U$ such that $\nabla\psi=0$ on $\partial\tilde{U}$ and $\psi$ is not realizable in $\tilde{U}$.
+Proof.
+Suppose that there exists an $x\in U$ such that $\nabla\psi(x)\neq 0$, otherwise, $\psi$ is constant and thus realizable in $U$.
+Then, by Lemma 1.3, $\psi$ is realizable by some $F$ in $B_{\epsilon}(x)$ for some small $\epsilon.$ Now let $U_{0}$ be the largest connected set containing $B_{\epsilon}(x)$, in which $\psi$ is realizable.
+By analyticity of $\psi$, $\psi$ is relizable by $F$ in $U_{0}$.
+Since $\psi$ is not realizable in $U$, $U_{0}$ must be a proper subset of $U$.
+We claim that $U_{0}$ cannot be simply connected.
+Indeed, if $U_{0}$ were simply connected, then $\partial U_{0}$ is a loop of critical points of $\psi$ and that $F$ is not analytic at $c=\psi(\partial{U}_{0})$.
+It follows from Proposition 4.9 that $\psi$ must be radially symmetric.
+This leads to a contradiction.
+We now take $\tilde{U}$ to be a simply-connected component of the set $\Omega\setminus U_{0}$.
+$\tilde{U}$ would be a simply connected domain, and the boundary $\tilde{\Gamma}$ is a piesewise analytic curve.
+Moreover, assume $\psi$ obtain value $\tilde{c}$ at $\tilde{\Gamma}$, then the $F$ ($F$ is previously defined in the range of $\psi$ in $\overline{U_{0}}$ and we have $\Delta\psi=F(\psi)$ in $\overline{U_{0}}$.) is not analytic at the value $\tilde{c}$.
+Now, suppose $\psi$ were realizable in $\tilde{U}$, say by $G$.
+We claim that $G$ is not analytic at $\tilde{c}$.
+Otherwise, let $\tilde{G}$ be the analytic extension of $G$.
+As $\Delta\psi-\tilde{G}(\psi)$ is an analytic function in a tubular neighborhood $\mathcal{N}$ of $\tilde{\Gamma}$ and $\Delta\psi-\tilde{G}(\psi)=0$ in $\tilde{U}$, we have $\Delta\psi-\tilde{G}(\psi)=0$ in $\mathcal{N}$.
+This, in particular, implies that $F$ is analytic at $\tilde{c}$, which leads to a contradiction.
+Since $G$ is not analytic at $\tilde{c}$, we could apply Proposition 4.9 to $\tilde{U}$ and conclude that $\psi$ is radially symmetric.
+Therefore, $\psi$ is not realizable in $\tilde{U}$.
+Clearly, by definition, $\tilde{U}\varsubsetneq U$.
+This completes the proof of the lemma.
+∎
+Based on Lemma 4.10, we give the proof of Theorem 1.1.
+Proof of Theorem 1.1.
+Suppose $\psi$ is neither radially symmetric nor realizable in $\Omega_{0}:=\Omega$, applying Lemma 4.10 iteratively, we obtain a nested family of simply connected domains $\{\Omega_{n}\}_{n=1}^{\infty}$ such that for $n\geq 1$, $\Omega_{n}\subsetneq\Omega_{n-1}$, $\psi$ is not realizable in $\Omega_{n}$, and $\nabla\psi=0$ on $\partial\Omega_{n}$.
+This, in particular, implies that there is an infinite family of nested loops of critical points for the analytic function $\psi$.
+It follows from Proposition 2.3 that $\psi$ is a constant function.
+This contradiction implies that $\psi$ must be radially symmetric or realizable in $\Omega$.
+∎
+4.3. Counterexample
 It is important to note that the rigidity in Theorem 1.1 is optimal in terms of regularity in bounded domains.
-First, the results of [18] show that Theorem 1.1 does not hold for flows that have finite smoothness even for the weaker assumption of non-locally-radial flows.
+First, the results of [23] show that Theorem 1.1 does not hold for flows that have finite smoothness, even for the weaker assumption of non-locally-radial flows.
 Secondly, Theorem 1.1 does not hold in the case where the compact domain $\Omega$ is not simply connected, as the following result shows.
-Theorem 4.12.
+Proposition 4.11.
 Fix an analytic Jordan curve $\gamma\subset\mathbb{R}^{2}$.
-Then there exists a tubular neighborhood $\Omega$ of $\gamma$ and a stream function $\psi\in C^{\omega}(\bar{\Omega})$ constant on $\partial\Omega$ solving
-$\begin{cases}\Delta\psi=2+\psi^{\frac{5}{2}}\text{ in }\Omega^{interior}\\ \Delta\psi=2-\psi^{\frac{5}{2}}\text{ in }\Omega^{exterior}\end{cases},\text{ and }\psi=\frac{\partial\psi}{\partial n}=0\text{ on }\gamma.$
-Thus, any non-circular $\gamma$ gives rise to a steady Euler flow $\psi$ that is analytic in a neighborhood of $\gamma$ that is neither radial nor a solution of a semilinear elliptic equation of the form (1.4).
-The proof of Theorem 4.12 is inspired by the classical proof the Cauchy-Kovalevskaya Theorem and a careful application of the so-called majorization method to a well chosen majorization system.
-A detailed proof is given in the thesis of the second named author [26].
-5 Conclusion and Future Directions
-We gather here a few interesting directions and problems that can be considered, in view of Theorem 1.1 and its proof.
+Then there exist a tubular neighborhood $\Omega:=\Omega^{int}\cup\gamma\cup\Omega^{ext}$ of $\gamma$ and a stream function $\psi\in C^{\omega}(\bar{\Omega})$, which is constant on $\partial\Omega$ and solves the following problem.
+[…]
+Thus, any non-circular $\gamma$ gives rise to a steady Euler flow $\psi$, that is analytic in a neighborhood of $\gamma$, and that is neither radial nor a solution of a semilinear elliptic equation of the form (1.4).
+In order to prove Proposition 4.11, we construct an analytic function $\psi$ in a tubular neighborhood $\overline{\Omega^{int}\cup\Omega^{ext}}$ of $\gamma$ satisfying
+$\Delta\psi=2+\psi^{\frac{5}{2}}\quad\text{in }\Omega^{int}.$
+The construction of such an analytic function is inspired by the classical proof of the Cauchy–Kovalevskaya Theorem, where we apply the so-called majorization method to a carefully chosen majorizing system, so that the existence of the analytic solution is established.
+Once such an analytic function $\psi$ is obtained, we can verify that $\{\psi,\Delta\psi\}=0$ in $\Omega^{ext}$. Analogous to the proof of Lemma 4.4, we can then prove
+$\Delta\psi=2-\psi^{\frac{5}{2}}\quad\text{in }\Omega^{ext}.$
+A detailed proof can be found in Sections 2.4.3–2.4.5 of the thesis of the second-named author [31].
+5. Conclusion and Future Directions
+We gather here a few interesting directions and problems that can be considered in view of Theorem 1.1 and its proof.
 We have already mentioned that replacing the analyticity assumption with some other less stringent assumption would be desirable.
 We mention some other interesting problems below.
-5.1 Local and Global Questions
+5.1. Local and Global Questions
 Given Theorem 1.1 and the extremely large literature on semilinear elliptic equations, an important problem is to study the local and global properties of the set of Euler steady states.
 As mentioned in the introduction, the invertibility properties of the linear Schrödinger operator
-$\mathcal{L}=\Delta-F_{*}^{\prime}(\psi_{*})$
+$\mathcal{L}=\Delta-F^{\prime}(\psi_{*})$
 play an important role in the local structure of steady states (see [12, 13, 15, 16, 36]).
 One consequence of Theorem 1.1 and the proof is that we can always define $\mathcal{L}$ as a (unbounded) self-adjoint operator on $L^{2}$.
-This gives hope to have a general local understanding of the manifold of steady states, and then possibly a global one.
+This gives hope for a general local understanding of the manifold of steady states and then possibly a global one.
 Most of the previous results on the local structure of steady states have assumed the invertibility of $\mathcal{L}$.
-It is a very interesting problem to study the general case where $\mathcal{L}$ has a kernel.
-From the classical Fredholm theory, we know the kernel is finite dimensional.
-When the kernel of $\mathcal{L}$ is one-dimensional, it is still possible to give a quite clear picture of the local structure of steady states (though this has not yet been spelled out in the literature).
-Once the kernel of $\mathcal{L}$ is two-dimensional or higher, the local structure of steady states may be quite complicated [11].
-It is a very interesting problem to study this in general and this is related to the (im)possibility of the existence of isolated states [14].
+It is an interesting problem to study the general case where $\mathcal{L}$ has a kernel.
+From the classical Fredholm theory, we know that the kernel is finite-dimensional.
+Once $\mathcal{L}$ has a kernel, the local structure of steady states may be quite complicated [15].
+It is an interesting problem to study this in general, and this is related to the (im)possibility of the existence of isolated states [20].
 If we had a good understanding of the local structure of steady states in the general case, we could hope to patch together this information to get more global information about the structure of the set of analytic Euler steady states.
-It is possible that some of these issues could be simpler to study on generic domains [42].
-For example, on generic domains it may even be possible to establish some kind of global correspondence between $\psi$ and $F$ in the analytic class [9].
-In general, it would be very interesting to have results related to the connectedness, path-connectedness, and smoothness of the manifold of steady states [12].
-5.2 Rigidity questions on other domains
+Some of these issues could be simpler to study in generic domains [49].
+In general, it would be interesting to have results related to the connectedness, path-connectedness, and smoothness of the manifold of steady states (see [16, 17]).
+5.2. Rigidity Questions on Other Domains
 A first obvious question is whether similar results can be established in different domains.
-We have already shown that it does not hold on some annular domains, though it is not clear whether a modification holds on $\mathbb{T}\times[0,1],$ for example.
-A very interesting problem is to determine whether a version of Theorem 1.1 can hold on $\mathbb{T}^{2}.$ In that case, the question would be whether all analytic steady states are necessarily either a shear flow or a solution to a semilinear elliptic equation.
+We have already shown that it does not hold in some annular domains, although it is not clear whether a modification holds on $\mathbb{T}\times[0,1],$ for example.
+An interesting problem is to determine whether a version of Theorem 1.1 can hold on $\mathbb{T}^{2}.$ In that case, the question would be whether all analytic steady states are necessarily either a shear flow or a solution to a semilinear elliptic equation.
 Acknowledgements
-T.M.E.
+T.
+M.
+E.
 acknowledges partial funding from the NSF DMS-2043024, the Alfred P.
@@
 Kiselev.
-Y.H.
+Y.
+H.
 acknowledges partial funding from NSF DMS-2043024.
@@
 is partially supported by NSFC grant 12250710674, and Program of Shanghai Academic Research Leader 22XD1421400.
-Appendix A Hopf Lemmas for elliptic equations with singular coefficents
-In this section, we list some known facts about the maximum principle for the solutions to the elliptic equation $-\Delta h+c(x)h=0$ in a Lipschitz domain $\Omega$.
-Here $c(x)$ could be unbounded, but it satisfies the following uniform upperbound $c^{-}(x)\leq\frac{M}{dist(x,\partial\Omega)}$.
+The authors thank the referees for many helpful comments that significantly improved the presentation of this paper.
+Appendix A Hopf Lemmas for Elliptic Equations with Singular Coefficients
+In this section, we list some known facts about the maximum principle for solutions to the elliptic equation $\Delta h+c(x)h=0$ in a Lipschitz domain $\Omega$.
+Here $c(x)$ could be unbounded but satisfies the following uniform upper bound $c^{-}(x)\leq\frac{M}{dist(x,\partial\Omega)}$.
 The results we present here are crucial in our proof of Proposition 4.3.
@@
 We consider a non-negative smooth solution $h$ to the following elliptic equation in the ball $B_{r}(0,r):$
-$-\Delta h+c(x)h=0.$ (A.1)
+(A.1) $\Delta h+c(x)h=0.$
 Now, if we further assume $h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial h}{\partial x_{2}}(0)>0$.
@@
 Let $B_{r}^{+}:=B_{r}(0,-r)\cap\{x_{1}\geq 0\}$, we consider a non-negative smooth solution $h$ to the following elliptic equation in $B_{r}^{+}:$
-$-\Delta h+c(x)h=0\text{ with $c^{-}(x)\leq\frac{M}{dist(x,0)}$.}$ (A.2)
-Now if we further assume $\nabla h(0)=h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial^{2}h}{\partial\eta^{2}}(0)>0$, for any direction $\eta=(\eta_{1},\eta_{2})$ satisfying $\eta_{1}>0$, $\eta_{2}<0$ in $B_{r}^{+}$.
+(A.2) $\Delta h+c(x)h=0\text{ with $c^{-}(x)\leq\frac{M}{dist(x,0)}$.}$
+Now, if we further assume $\nabla h(0)=h(0)=0$, there exists a $r_{0}>0$ depending on $M$, such that if $r<r_{0}(M)$, we have either $h=0$ or $\frac{\partial^{2}h}{\partial\eta^{2}}(0)>0$, for any direction $\eta=(\eta_{1},\eta_{2})$ satisfying $\eta_{1}>0$, $\eta_{2}<0$ in $B_{r}^{+}$.
+Appendix B On the Rigidity of Steady States with Single-Signed Vorticity
+We now proceed to make precise statements in Remark 1.5.
+Proposition B.1.
+Let $\Omega$ be a simply connected, bounded, and convex domain, and $\psi$ be the stream function of a steady Euler flow with non-positive vorticity in $\Omega$.
+If we further assume that $\Omega$ is even symmetric in the direction $e:=(1,0)$, and $\psi$ is analytic, then $\psi$ is even symmetric in the direction $e$.
+The full proof of this theorem follows inverbatim from the proof of Theorem 1.1.
+We explain the key ideas here.
+In the case where $\psi$ is nonradial, by Theorem 1.1, there exists an $F$ such that $\Delta\psi=F(\psi)$ in $\Omega$.
+Without loss of generality, we assume $\psi=0$ in $\partial\Omega$.
+As $\Delta\psi$ is non-positive, unless $\psi=0$, we have $\psi>0$ in $\Omega$ and $F$ is analytic near 0.
+Without loss of generality, $max_{\Omega}\psi=1$, then $F$ is analytic in $(0,1)$ and admits a Puiseux expansion near $1$ as Lemma 4.4.
+Analogously to the proof of Lemma 4.7, we can show $\psi$ is even symmetric in direction $e$.
+Remark B.1.
+An interesting question is to extend the study in Proposition B.1 when the domain is non-convex.
+More precisely, in bounded simply connected domains with mirror symmetry, determine whether analytic steady flows with single-signed vorticity should enjoy mirror symmetry.